### v1

# Characters of symmetric groups: sharp bounds on virtual degrees and the Witten zeta function

Abstract
We prove sharp bounds on the virtual degrees introduced by Larsen and Shalev. This leads to improved bounds on character of symmetric groups. We then sharpen bounds of Liebeck and Shalev concerning the Witten zeta function. As an application, we characterize the fixed-point free conjugacy classes whose associated random walk mixes in 2 steps.
Résumé
Nous prouvons des bornes précises sur les degrés virtuels introduits par Larsen et Shalev. Cela induit de meilleures bornes sur les caractères du groupe symétrique. Dans un second temps, nous améliorons certaines bornes de Liebeck et Shalev sur la fonction zeta de Witten. En guise d’application, nous caractérisons les classes de conjugaison sans point fixe dont la marche aléatoire associée est mélangée au temps 2.
Resumo
Ni pruvas precizajn barojn je la virtualaj gradoj enkondukitaj de Larsen kaj Shalev. Tio induktas plibonigitajn barojn je karakteroj de la simetria grupo. Ni sekve pliakrigas barojn de Liebeck kaj Shalev pri la zeto-funkcio de Witten. Kiel aplikaĵo, ni karakterizas la senfikspunktajn konjugklasojn kies asociata hazarda promenado miksiĝas post 2 paŝoj.
1 Introduction
1.1 Context
Representation theory has been used to solve various problems in different areas of mathematics. A striking example is the pioneering work to Diaconis and Shahshahani on a random process called the random transposition shuffle [DS81]. They used character estimates to prove a sharp phase transition - now called cutoff - concerning the minimum number of random transpositions whose product is an almost uniform random permutation. This was the starting point of the field of mixing times.
Other techniques to study mixing times were developed in the following years, notably by Aldous and Diaconis [Ald83, AD86], and mixing properties of emblematic card shufflings were precisely understood, also for small decks of cards [BD92]. We refer to [LP17] for an introduction to mixing times and to [DF23] for the mathematics of card shuffling.
A natural way to generalize Diaconis and Shahshahani’s random transposition shuffle on the symmetric group ${\mathfrak{S}}_{n}$ is to replace the conjugacy class of transpositions - from which we pick uniform elements - by another conjugacy class. In specific cases, a similar cutoff phenomenon was proved ; for $k$-cycles (where $k>n/2$) in [LP02] ; for $k$-cycles (where $k$ is fixed) in [BSZ11] ; for $k$-cycles (where $k=o(n)$) in [Hou16] ; and for all conjugacy classes whose support have size $o(n)$ in [BŞ19]. Most results rely on representation theory, but interestingly, the proof of [BSZ11] follows the cycle structure of the permutation obtained after several multiplications via probabilistic arguments, and that of [BŞ19], which is currently the best result for conjugacy classes with small support, relies on a curvature argument.
For conjugacy classes with cycle type $[r^{n/r}]$, Lulov [Lul96] proved that the mixing time is $3$ if $r=2$ and $2$ if $r\geq 3$. Lulov and Pak then conjectured in [LP02, Conjecture 4.1] that the mixing time is at most 3 for all fixed-point free conjugacy classes (i.e. without fixed point), and Larsen and Shalev [LS08] proved that conjecture.
In the last years, results were obtained concerning the cutoff profile of such processes, that is, what happens within the cutoff window (when we zoom in, in the phase transition). The cutoff profile for the random transposition shuffle was found by the first named author in [Tey20], and this was later generalized to $k$-cycles (for $k=o(n)$) by Nestoridi and Olesker-Taylor [NOT21]. Recently, another proof for the profile for transpositions was obtained by Jain and Sawhney [JS24], using a hybrid technique that combines probabilistic and representation theoretic arguments.
Motivated among other problems by the study of mixing times of Markov chains, uniform bounds on characters were also established and sharpened over the last decades. Roichman [Roi96] proved bounds that led to the correct order of the mixing time of conjugacy class walks on $\mathfrak{S}_{n}$ that have a number of fixed points of order $n$. This was later strengthened by Müller and Schlage-Puchta [MSP07], who obtained results that are uniform over all conjugacy classes.
Sharp bounds for characters of permutations with few fixed points were proved in the landmark paper [LS08] of Larsen and Shalev. They proved several conjectures, among which the Lulov–Pak conjecture as mentioned above.
The bounds that we obtain build on this paper ; see Section 1.2 for a more detailed description of the character bounds of Larsen and Shalev.
An important tool to estimate characters of ${\mathfrak{S}}_{n}$ is the Murnaghan–Nakayama rule (see e.g. Theorem 3.10 in [Mé17]), which is a key element in the proofs of [Roi96, MSP07, LS08]. On the other hand, other techniques have been proved to be more powerful in some regimes. Féray and Śniady [FS11] found uniform bounds on characters that are especially good for diagrams that do not have a long first line or column, using a reformulation of Stanley’s conjectured character formula. More precisely, [Fé10] proves Conjecture 3 from [Sta06], and the reformulation is Theorem 2 in [FS11].
In another direction, Lifschitz and Marmor [LM23] recently used a new technique called hypercontractivity to prove character bounds. They obtain uniform bounds on characters for permutations having at most a given number of cycles (instead of considering the full cycle structure, which was done in the previously mentioned works). This allows them to consider permutations having up to $n/(\ln n)^{O(1)}$ cycles and therefore cover some new cases.
Bounds on characters are very important in the more general context of finite groups, and were notably previously used to prove a conjecture due to Ore [LOST10]. General bounds on characters of finite classical groups were recently obtained by Guralnick, Larsen and Tiep [GLT20, GLT24], and uniform bounds on characters for all finite quasisimple groups of Lie type were established by Larsen and Tiep [LT24].
A canonical way to convert character bounds for a group $G$ is via the Witten zeta function. It is defined as $\zeta(s)=\sum_{\alpha}(d_{\alpha})^{-s}$ for $s>0$, where the sum is over all irreducible representations $\alpha$ of $G$, and $d_{\alpha}$ denotes the dimension of $\alpha$. Such a zeta function appeared in the computation of volumes of moduli spaces in the work of Witten [Wit91] (see Equation (4.72) there). Liebeck and Shalev proved bounds for the symmetric group [LS04] - that were used by Larsen and Shalev [LS08] in their applications to mixing times and covering numbers - and for Fuchsian groups [LS05].
We refer to [LS08], [LT24], and the survey article [Lie17], for more applications of characters bounds.
We will present further applications in a follow-up paper [TT24]. We are also developing a new technique to find character bounds, that we will present in [OTTT24].
In the whole paper, $\mathfrak{S}_{n}$ will denote the symmetric group of order $n$, and $\widehat{\mathfrak{S}_{n}}$ the set of its irreducible representations. For a permutation $\sigma$ and $i\in\{1,\ldots,n\}$, $f_{i}(\sigma)$ denotes the number of cycles of length $i$ in $\sigma$. By extension, ${\mathcal{C}}$ being a conjugacy class, we will denote by $f_{i}({\mathcal{C}})$ the number of cycles of length $i$ of any permutation $\sigma\in{\mathcal{C}}$. For convenience, we will use $\lambda$ to denote at the same time a representation, the associated integer partition, and the associated Young diagram (see Section 2.1 for definitions). If $E$ is a finite set, we denote by $\operatorname{\mathrm{Unif}}_{E}$ the uniform probability measure on $E$.
1.2 The Larsen–Shalev character bounds
We describe here the results of Larsen and Shalev [LS08], whose improvement is the main purpose of the paper. We assume familiarity with the representation theory of symmetric groups and will follow the notation from the textbook [Mé17]: we denote by $d_{\lambda}$ the dimension of a representation $\lambda$, $\operatorname{\mathrm{ch}}^{\lambda}(\sigma)$ the character of $\lambda$ evaluated at a permutation $\sigma\in\mathfrak{S}_{n}$, and $\chi^{\lambda}(\sigma)=\frac{\operatorname{\mathrm{ch}}^{\lambda}(\sigma)}{d_{\lambda}}$ the associated normalized character.
Larsen and Shalev obtained character bounds by introducing and studying new objects. One of them is called the virtual degree $D(\lambda)$ of a representation $\lambda$. It is defined as
$D(\lambda):=\frac{(n-1)!}{\prod_{i}a_{i}!b_{i}!},$ (1.1)
where $a_{i}=\lambda_{i}-i$ and $b_{i}=\lambda^{\prime}_{i}-i$. Here $\lambda_{i}$ (resp. $\lambda^{\prime}_{i}$) denotes the size of the $i$-th row (resp. $i$-th column) of the Young diagram associated to $\lambda$ (see Section 2.1 for precise definitions). The virtual degree is a convenient substitute for the dimension $d_{\lambda}$, which allows for easier computations. See Section 2.3 for more details.
Another object introduced in [LS08] is the orbit growth exponent $E(\sigma)$. Let $n\geq 1$ and $\sigma\in\mathfrak{S}_{n}$. Denote by $f_{i}(\sigma)$ the number of cycles of length $i\geq 1$ of $\sigma$, and for $k\geq 1$, set $\Sigma_{k}=\Sigma_{k}(\sigma):=\sum_{1\leq i\leq k}if_{i}(\sigma)$. We define the orbit growth sequence $(e_{i})_{i\geq 1}$ by $n^{e_{1}+\ldots+e_{i}}=\max\left(\Sigma_{i},1\right)$ for all $i\geq 1$, and set
$E(\sigma)=\sum_{i\geq 1}\frac{e_{i}}{i}.$ (1.2)
A key intermediate result that Larsen and Shalev proved is the following, which we will refer to as a theorem.
Theorem 1.1 ([LS08]).
For any $n\geq 1$, $\sigma\in\mathfrak{S}_{n}$ and $\lambda\in\widehat{\mathfrak{S}_{n}}$, we have the following character bound:
$\left|\operatorname{\mathrm{ch}}^{\lambda}(\sigma)\right|\leq D(\lambda)^{E(\sigma)}.$ (1.3)
This is proved via an elegant induction in the proof of their main theorem (for reference, the induction hypothesis is [LS08, Eq. (17)]), where the orbit growth exponent $E(\sigma)$ naturally appears. They also showed that $D(\lambda)$ cannot be much larger than $d_{\lambda}$.
Theorem 1.2 ([LS08], Theorem 2.2).
As $|\lambda|\to\infty$, we have
$D(\lambda)\leq d_{\lambda}^{1+o(1)}.$ (1.4)
Combining Theorem 1.1, Theorem 1.2 and the fact that $E(\sigma)\leq 1$ for all $\sigma$, they obtain the following character bound.
Theorem 1.3 ([LS08], Theorem 1.1 (a)).
As $n\to\infty$, we have uniformly over all $\sigma\in\mathfrak{S}_{n}$ and $\lambda\in\widehat{\mathfrak{S}_{n}}$,
$\left|\operatorname{\mathrm{ch}}^{\lambda}(\sigma)\right|\leq d_{\lambda}^{E(\sigma)+o(1)}.$ (1.5)
The orbit growth exponent $E(\sigma)$ defined in (1.2) looks complicated at first glance, but it is actually simple to compute and appears to lead to the best possible character bounds in different regimes such as in (1.6) below.
Let us give a few important examples of values of $E(\sigma)$. We always denote by $f_{1}$ the number of fixed points of a permutation $\sigma$ and set $f=\max(f_{1},1)$.
Example 1.4.
  1. (a)
    Assume that $\sigma\sim[2^{n/2}]$, then $e_{2}=1$ and $e_{i}=0$ for $i\neq 2$, so $E(\sigma)=1/2$.
  2. (b)
    Assume that $\sigma$ has no fixed point, then $e_{2}+e_{3}+...=1$ so $E(\sigma)\leq 1/2$.
  3. (c)
    Assume that $\sigma$ has at least one fixed point, then $e_{1}=\frac{\ln f}{\ln n}$ so $E(\sigma)\leq e_{1}+\frac{1}{2}(1-e_{1})=\frac{1+e_{1}}{2}=\frac{1}{2}+\frac{\ln f}{2\ln n}$, that is $E(\sigma)-1=-\frac{1}{2}\frac{\ln n/f}{\ln n}$
  4. (d)
    Assume that $\sigma\sim[k,1^{n-k}]$, then $E(\sigma)=1/n$ if $k=n$ (i.e. if $\sigma$ is an $n$-cycle) and if $f_{1}\geq 1$, $E(\sigma)=e_{1}+\frac{e_{k}}{k}=e_{1}+\frac{1-e_{1}}{k}=\frac{\ln f}{\ln n}+\frac{\ln(n/f)}{k\ln n}$.
Combining Theorem 1.3 with Example 1.4 (b), Larsen and Shalev obtained the following bound as $n\to\infty$, uniform over fixed-point free permutations $\sigma\in\mathfrak{S}_{n}$ and irreducible representations $\lambda\in\widehat{\mathfrak{S}_{n}}$,
$\left|\operatorname{\mathrm{ch}}^{\lambda}(\sigma)\right|\leq d_{\lambda}^{1/2+o(1)}.$ (1.6)
This allowed them to prove the abovementioned conjecture of Lulov and Pak, and even extend it to the mixing time of conjugacy classes that have a small number of fixed points (see [LS08, Theorem 1.8].
1.3 Main results
We present here our main contributions: the main result, Theorem 1.5, provides uniform bounds on virtual degrees and allows us to obtain refined character bounds (Theorem 1.6). Proposition 1.7 provides bounds on the Witten zeta function, which we apply together with Theorem 1.6 to characterize fixed-point free conjugacy classes that mix in 2 steps (Theorem 1.8).
Our main result is the following technical improvement to Theorem 1.2.
Theorem 1.5.
There exists a universal constant $C$ such that, for every $n\geq 2$ and any integer partition $\lambda\vdash n$, we have
$D(\lambda)\leq d_{\lambda}^{1+\frac{C}{\ln n}}.$ (1.7)
We will show in Section 4.4 that Theorem 1.5 is sharp for different shapes of diagrams.
Plugging this into Theorem 1.1, we get the following character bound.
Theorem 1.6.
For every $n\geq 2$, any permutation $\sigma\in\mathfrak{S}_{n}$ and any integer partition $\lambda\vdash n$, we have
$\left|\operatorname{\mathrm{ch}}^{\lambda}(\sigma)\right|\leq d_{\lambda}^{\left(1+\frac{C}{\ln n}\right)E(\sigma)},$ (1.8)
where $C$ is the universal constant from Theorem 1.5.
Remark 1.1.
Larsen and Shalev [LS08, Theorem 1.3] obtained the bound $\frac{|\operatorname{\mathrm{ch}}^{\lambda}(\sigma)|}{d_{\lambda}}\leq d_{\lambda}^{-\frac{1}{2}\frac{\ln n/f}{\ln n}+o(1)}$, where $f=\max(1,f_{1}(\sigma))$. Theorem 1.6 allows us to improve it to $\frac{|\operatorname{\mathrm{ch}}^{\lambda}(\sigma)|}{d_{\lambda}}\leq d_{\lambda}^{-\frac{1}{2}\frac{\ln n/f}{\ln n}+O\left(\frac{1}{\ln n}\right)}$. In particular, when $f=o(n)$, our result provides the bound
$\frac{|\operatorname{\mathrm{ch}}^{\lambda}(\sigma)|}{d_{\lambda}}\leq d_{\lambda}^{-\left(\frac{1}{2}+o(1)\right)\frac{\ln n/f}{\ln n}}.$ (1.9)
We are working on extending such a bound to all permutations.
On the other hand, we define the Witten zeta function by, for $n\geq 1$, $A\subset\widehat{\mathfrak{S}_{n}}$, and $s\geq 0$,
$\zeta_{n}(A,s):=\sum_{\lambda\in A}\frac{1}{\left(d_{\lambda}\right)^{s}}.$ (1.10)
Note that compared to the usual definition where the sum is over all irreducible representations, we add the subset of representations $A$ on which we sum as a parameter of the function $\zeta_{n}$.
In [LS04], Liebeck and Shalev proved that $\zeta_{n}(\widehat{\mathfrak{S}_{n}}^{**},s)=O(n^{-s})$, if $s>0$ is fixed and $n\to\infty$, where $\widehat{\mathfrak{S}_{n}}^{**}=\widehat{\mathfrak{S}_{n}}\backslash\left\{[n],[1^{n}]\right\}$. We improve their result to allow the argument $s$ to tend to 0.
Proposition 1.7.
Let $(s_{n})$ be a sequence of positive real numbers. We have
[…]
In Section 5, we will also prove variants of Proposition 1.7, for other subsets $A\subset\widehat{{\mathfrak{S}}_{n}}$, and when $s_{n}$ is of order $\frac{1}{\ln n}$.
As an application of our bounds on characters and on the Witten zeta function, we are able to characterize which fixed-point free conjugacy classes mix in 2 steps.
Theorem 1.8.
For each $n\geq 2$ let ${\mathcal{C}}^{(n)}$ be a conjugacy class of ${\mathfrak{S}}_{n}$, which is fixed-point free (i.e. $f_{1}({\mathcal{C}}^{(n)})=0$). Recall that $f_{2}({\mathcal{C}}^{(n)})$ denotes the number of transpositions of a permutation $\sigma\in{\mathcal{C}}^{(n)}$. Then we have
$\operatorname{d_{TV}}\left(\operatorname{\mathrm{Unif}}_{{\mathcal{C}}^{(n)}}^{*2},\operatorname{\mathrm{Unif}}_{{\mathfrak{A}}_{n}}\right)\xrightarrow[n\to\infty]{}0\quad\quad\text{ if and only if }\quad\quad f_{2}({\mathcal{C}}^{(n)})=o(n).$ (1.12)
We believe that the only parts of the cycle structure that affect mixing times are the number of fixed points and the number of transpositions. We make the following conjecture.
Conjecture 1.9.
For each $n\geq 2$ let ${\mathcal{C}}^{(n)}$ be a conjugacy class of ${\mathfrak{S}}_{n}$. Then
$\operatorname{d_{TV}}\left(\operatorname{\mathrm{Unif}}_{{\mathcal{C}}^{(n)}}^{*2},\operatorname{\mathrm{Unif}}_{{\mathfrak{A}}_{n}}\right)\xrightarrow[n\to\infty]{}0\quad\text{ if and only if }\quad[f_{1}=o(\sqrt{n})\text{ and }f_{2}=o(n)],$ (1.13)
where $f_{1}=f_{1}({\mathcal{C}}^{(n)})$ and $f_{2}=f_{2}({\mathcal{C}}^{(n)})$.
1.4 Structure of the paper
Section 2 is devoted to preliminaries on Young diagrams and some elementary results. We recall there in particular the celebrated hook length formula.
In Section 3.1, we introduce the sliced hook products. This generalizes the idea of replacing the whole hook product of a diagram by a simpler product to any set partition of a diagram. In Section 3.2, we study sliced hook products and obtain precise approximations of (classical) hook products. A key result is the bound $d_{\lambda}\geq\binom{n}{|c|}d_{s}d_{c}e^{6\sqrt{|c|}}$ presented in Proposition 3.8 (b), which improves on [LS08, Lemma 2.1]. Here, $s$ denotes the external hook of $\lambda$ and $c=\lambda\backslash s$ is the center of $\lambda$ (as drawn in Section 2.3). In Section 3.3, we introduce the notion of augmented dimension $d_{\lambda}^{+}$ and consider some of its useful properties.
In Section 4 we use the results of Section 3.2 to prove Theorem 1.5 in two steps. We first show in Section 4.2 that the virtual degree $D(\lambda)$ and the augmented dimension $d_{\lambda}^{+}$ are close to each other. We then prove in Section 4.3, by induction, slicing the external hooks of diagrams $\lambda$, that $d_{\lambda}^{+}=d_{\lambda}^{1+O(1/\ln|\lambda|)}$. Finally, in Section 4.4 we provide examples that show that Theorem 1.5 is sharp up to the value of the constant.
In Section 5, we improve the bounds from Liebeck and Shalev [LS04] on the Witten zeta function and prove Proposition 1.7.
Finally, in Section 6, we apply the previous results and characterize which fixed-point free conjugacy classes mix in 2 steps, proving Theorem 1.8.
2 Preliminaries
2.1 Young diagrams, representations and integer partitions
We say that $\lambda:=[\lambda_{1},\ldots,\lambda_{k}]$ is a partition of the integer $n$ if $\lambda_{i}\in\mathbb{Z}_{\geq 1}:=\{1,2,\ldots\}$ for all $1\leq i\leq k$, $\lambda_{i}\geq\lambda_{i+1}$ for all $i\leq k-1$, and $\sum_{i=1}^{k}\lambda_{i}=n$. We write $\lambda\vdash n$ if $\lambda$ is a partition of $n$. It turns out that integers partitions of $n$ are in bijection with irreducible representations of ${\mathfrak{S}}_{n}$, which makes it a good tool to study characters.
A common and useful way to code an integer partition (and thus a representation of ${\mathfrak{S}}_{n}$) is through Young diagrams. The Young diagram of shape $\lambda:=\left[\lambda_{1},\ldots,\lambda_{k}\right]$ (where $\lambda$ is a partition of an integer $n$) is a table of boxes, whose $i$-th line is made of $\lambda_{i}$ boxes, see Fig. 1 for an example.
[…]
Figure 1: The Young diagram coding the partition $[5,2,1]$ of the integer $8$. It has 5 boxes on the first line, 2 on the second, and 1 on the third.
We also denote by $\lambda$ this Young diagram, and say that $n$ is its size.
2.2 Hook lengths, hook products and the hook-length formula
Let $\lambda$ be a Young diagram. If $u\in\lambda$, the hook of $u$ in $\lambda$ is the set of boxes on the right or above $u$, including $u$. We call hook length the size of this set and denote it by $H(\lambda,u)$. See Fig. 2 for an example.
[…]
*(orange)
[…]
*(pink!60)
[…]
*(pink!60) *(pink!60) *(pink!60) *(pink!60)
[…]
Figure 2: Left: the Young diagram associated to $\lambda=[7,5,4,1]$, with a box $u$ colored in orange. Right: the hook associated to $u$ is in pink. Its length is $H(\lambda,u)=5$.
Let us now consider a subset of boxes $E\subset\lambda$. We define
$H(\lambda,E):=\prod_{u\in E}H(\lambda,u)$ (2.1)
the hook-product of $E$ in $\lambda$. See Fig. 3 for an example.
[…]
*(orange)
[…]
*(orange) *(orange)
[…]
Figure 3: The Young diagram coding $\lambda=[7,5,4,1]$, with the set $E=\left\{(1,1),(3,1),(6,1),(1,2),(2,2),(3,3)\right\}$ in orange. Here, $H(\lambda,E)=10\cdot 7\cdot 2\cdot 7\cdot 5\cdot 2=9800$.
We can now state the hook-length formula. Let $\lambda$ be a Young diagram and $n:=|\lambda|$ its number of boxes. A standard tableau of $\lambda$ is a filling of $\lambda$ with the numbers from 1 to $n$ such that the numbers are increasing on each line and column. We denote by $\operatorname{ST}(\lambda)$ the set of standard tableaux of $\lambda$, and $d_{\lambda}:=|\operatorname{ST}(\lambda)|$, see Fig. 4. It is well-known that $d_{\lambda}$ is the dimension of the irreducible representation of the symmetric group associated to $\lambda$.
[…]
& *(green!60)4 *(cyan!60)5
[…]
Figure 4: The diagram $\lambda=[3,2]$ and its $d_{\lambda}=|\operatorname{ST}(\lambda)|=5$ standard tableaux.
The hook-length formula was discovered in 1953 by Frame, Robinson, and Thrall [FRT54], and allows to compute the number of standard tableaux of a diagram looking only at its hook lengths: for any diagram $\lambda$ of size $n$, we have
$d_{\lambda}=\frac{n!}{H(\lambda,\lambda)}.$ (2.2)
Example 2.1.
Let us consider again the $\lambda=[3,2]$, which has size 5. We have $\frac{n!}{H(\lambda,\lambda)}=\frac{5!}{4\cdot 3\cdot 1\cdot 2\cdot 1}=\frac{120}{24}=5$. Therefore, by the hook length formula, we recover that $d_{\lambda}=5$.
2.3 Diagram notation
Let us define some notation that we will use for diagrams. To keep the notation light, we will usually use the same pieces of notation for both the diagrams and their lengths. For example we will say the first line of $\lambda$ is $\lambda_{1}$ and has length $\lambda_{1}$. If we want to emphasize that we are considering a diagram or a size, we will add brackets or absolute values. For example we can say that the first line $\left[\lambda_{1}\right]$ has length $\left|\lambda_{1}\right|$.
Definition 2.2 (Definition-notation).
Let $\lambda$ be a Young diagram.
  • We denote by $\lambda_{i}$ the $i$-th line of $\lambda$ and by $\lambda_{i}^{\prime}$ its $i$-th column.
  • We denote by $\delta_{i}$ the $i$-th diagonal box of $\lambda$, that is, $\delta_{i}=\lambda_{i}\cap\lambda^{\prime}_{i}$.
  • We denote by $\delta=\delta(\lambda)$ the diagonal of $\lambda$, that is, $\delta=\cup_{i}\delta_{i}$.
  • We denote by $r$ the truncated diagram (removing the first line of $\lambda$, $s=s(\lambda):=\lambda_{1}\cup\lambda_{1}^{\prime}$ the external hook of $\lambda$, and by $c=\lambda\backslash s$ its center (i.e. the diagram, except the first line and the first column).
  • We denote $\lambda_{\geq i}:=\cup_{j\geq i}\lambda_{j}$ and $\lambda_{\leq i}:=\cup_{j\leq i}\lambda_{j}$.
  • Similarly, we define $\lambda^{\prime}_{\geq i}:=\cup_{j\geq i}\lambda^{\prime}_{j}$, $\lambda^{\prime}_{\leq i}:=\cup_{j\leq i}\lambda^{\prime}_{j}$, and $\lambda^{\prime}_{i_{1}\to i_{2}}:=\cup_{i_{1}\leq j\leq i}\lambda^{\prime}_{j}$.
  • We denote by $\lambda_{a}^{i}=a_{i}$ and $\lambda_{b}^{i}=b_{i}$ the truncated $i$-th lines and columns of $\lambda$.
  • We denote by $\lambda^{i}$ the $i-th$ hook of $\lambda$, that is the boxes that are on the right of $\delta_{i}$ on $\lambda_{i}$ or above $\delta_{i}$ on $\lambda^{\prime}_{i}$. Formally we can write $\lambda^{i}=(\lambda_{i}\cap\lambda^{\prime}_{\geq i})\cup(\lambda_{\geq i}\cap\lambda^{\prime}_{i})$.
  • We define $\lambda^{\geq i}:=\cup_{j\geq i}\lambda^{j}$, $\lambda^{\leq i}:=\cup_{j\leq i}\lambda^{j}$.
  • We denote by $u_{i}$ the part of the $i$-th column which is above the first line. That is $u_{i}=\lambda_{i}^{\prime}\cap\lambda_{\geq 2}$.
  • We define $u_{\geq i}:=\cup_{j\geq i}u_{j}$, $u_{\leq i}:=\cup_{j\leq i}u_{j}$.
  • We denote by $\lambda_{\geq i,\geq j}$ the subdiagram of $\lambda$ whose boxes are on a line after the $i$-th and on a column after the $j$-th.
Fig. 5 illustrates these definitions on diagrams.
[…]
*(red!60)
[…]
*(red!60)*(red!60)*(red!60)
[…]
*(red!60)*(red!60)*(red!60)*(red!60)
[…]
*(red!60)
[…]
*(red!60)*(red!60)*(red!60)
[…]
[u_2]
[…]
*(red!60)
[…]
*(red!60)*(red!60)*(red!60)
[…]
*(orange!90)*(orange!90)*(orange!90) *(orange!90)*(orange!90)
[…]
*(orange!90)*(orange!90)*(orange!90) *(orange!90) *(orange!90) *(orange!90)
[…]
& *(orange!90)
[…]
*(orange!90)*(orange!90)*(orange!90)
[…]
*(orange!90)*(orange!90)*(orange!90)*(orange!90)*(orange!90)
[…]
Figure 5: Examples of Definition 2.2.
2.4 Some elementary results
We collect here standard or elementary results that we will us throughout the paper. We give proofs for completeness.
Lemma 2.3.
  1. (a)
    For any $n\geq 1$ and $\lambda\vdash n$ we have $d_{\lambda}\leq\sqrt{n!}$.
  2. (b)
    For $n\geq 6$, we have $n!\leq(n/2)^{n}$.
  3. (c)
    Let $n\geq 1$ and $\lambda\vdash n$. Denote by $s=s(\lambda)$ the external hook of $\lambda$. Then $d_{s}=\binom{s-1}{\lambda_{a}^{1}}$.
Proof.
  1. (a)
    A classical formula from representation theory is the following: for any finite group $G$ we have $|G|=\sum_{\rho\in\widehat{G}}d_{\rho}^{2}$, where $\widehat{G}$ is the set of all irreducible representations of $G$. This identity represents for example the dimensions on both sides of the Fourier isomorphism (see [Mé17, Section 1.3]). Hence, for any $\lambda\in\widehat{G}$ of a finite group $G$ we have
    $d_{\lambda}^{2}\leq\sum_{\rho\in\widehat{G}}d_{\rho}^{2}=|G|.$ (2.3)
    The result follows taking square roots on both sides, since for $n\geq 1$ we have $|\mathfrak{S}_{n}|=n!$.
  2. (b)
    We proceed by induction. The result holds for $n=6$ since $6!=720\leq 729=3^{6}=(6/2)^{6}$. Let now $n\geq 6$ and assume that $n!\leq(n/2)^{n}$. Then
    $\frac{(n+1)!}{((n+1)/2)^{n+1}}=\frac{n!}{(n/2)^{n}}(n+1)\frac{2}{n+1}\left(\frac{n}{n+1}\right)^{n}\leq 2\left(\frac{n}{n+1}\right)^{n}.$ (2.4)
    Since $\ln$ is concave, we have $\ln(1+x)\geq x\ln 2$ for $0\leq x\leq 1$, and therefore, we have for $n\geq 1$
    $2\left(\frac{n}{n+1}\right)^{n}=2e^{-n\ln(1+1/n)}\leq 2e^{-n\frac{\ln 2}{n}}=1,$ (2.5)
    concluding the proof.
  3. (c)
    Recall that $d_{s}$ is the number of standard tableaux of $s$. Observe that a standard tableau necessarily has 1 placed in the bottom-left corner. Furthermore, since $s$ is a hook, the rest of the standard diagram is determined by which numbers we chose to place on the first line, for which there are $\binom{s-1}{\lambda_{a}^{1}}$ possibilities.
3 Sliced hook products
We introduce here a new tool in the study of Young diagrams, which we call sliced hook products. This notion of hook product, which consists in cutting a diagram along its lines and columns, turns out to be suited for proofs by induction.
3.1 Definitions
Despite being elegant, the hook-length formula may be tricky to use, because of how hard it is to estimate hook products. The idea is to take into account only part of the hook lengths, making the resulting product easier to compute and to use.
We first extend the definition of hook lengths to any set of boxes. We represent boxes by the coordinates of their top right corner in the plane ${\mathbb{Z}}^{2}$. As such a box can be seen as an element of ${\mathbb{Z}}^{2}$ and a set of boxes is a subset $S\subset{\mathbb{Z}}^{2}$, see Fig. 6.
Definition 3.1.
Let $S\subset{\mathbb{Z}}^{2}$ be a finite set of boxes, and let $u=(x,y)\in{\mathbb{Z}}$ (not necessarily in $S$). We define the hook length of $u$ with respect to $S$ by
$H(S,u)=\left|\left\{(x^{\prime},y^{\prime})\in S{\;:\;}[x=x^{\prime}\text{ and}y\leq y^{\prime}]\text{ or }[y=y^{\prime}\text{ and }x\leq x^{\prime}]\right\}\right|.$ (3.1)
[…]
*(pink!60)
[…]
Figure 6: The hook of $u=(3,1)$ in the set of boxes $S=\left\{(1,1),(4,1),(5,1),(7,1),(3,2),(4,2),(1,3),(1,4),(3,4)\right\}$. In pink are the boxes in the hook of $u$, and we have $H(S,u)=5$.
Definition 3.2.
Let $\lambda$ be an integer partition and $P=\left\{\nu_{1},\nu_{2},...,\nu_{r}\right\}$ be a set partition of $\lambda$ (i.e. such that no $\nu_{i}$ is empty and $\sqcup_{i}\nu_{i}=\lambda$). The hook product sliced along $P$ is the product
$H^{*}(\lambda,\lambda)=H^{*P}(\lambda,\lambda):=\prod_{i=1}^{r}H(\nu_{i},\nu_{i}).$ (3.2)
If $u\in\nu_{i}$, we call the quantity $H(\nu_{i},u)$ the hook length of $u$ sliced along $P$, or sliced hook length.
More generally, if $E$ is a subset of $\lambda$, we define the following (partial) sliced hook product:
$H^{*P}(\lambda,E):=\prod_{i=1}^{r}H(\nu_{i},\nu_{i}\cap E).$ (3.3)
We will refer to “slicing” for the procedure consisting of replacing $H(\lambda,\lambda)$ by $H^{*}(\lambda,\lambda)$, as well as for the associated partition $P$.
Let us give some examples of slicings. We will each time provide an example where the $\nu_{i}$ are represented in different colors, and the boxes of the diagrams $\lambda$ are filled with their sliced hook lengths.
Definition 3.3.
Let $\lambda$ be a Young diagram.
  • We call $\lambda_{1}$-slicing (of $\lambda$) a slicing along $\left\{\lambda_{1},\lambda_{\geq 2}\right\}$, see Fig. 7.
  • We call $\lambda^{1}$-slicing (of $\lambda$) a slicing along $\left\{\lambda^{1},\lambda^{\geq 2}\right\}$, see Fig. 8.
  • We call $ab\delta$-slicing (of $\lambda$) a slicing along $P:=\left\{\lambda^{i}_{a}\right\}_{i}\sqcup\left\{\lambda^{i}_{b}\right\}_{i}\sqcup\left\{\delta_{i}\right\}_{i}$, see Fig. 9. We denote the (partial) $ab\delta$-sliced hook products by $H^{*ab\delta}(\lambda,\cdot)$.
[…]
&*(red!60)1
[…]
*(red!60)3*(red!60)1
[…]
*(red!60)4*(red!60)2
[…]
*(red!60)8*(red!60)6*(red!60)3*(red!60)1
[…]
Figure 7: The $\lambda_{1}$-slicing for $\lambda=[14,4,3,2,2,1]$. Numbers in the boxes corespond to the hook lengths after slicing.
[…]
&*(blue!60)1
[…]
*(blue!60)2*(red!60)1
[…]
*(blue!60)3*(red!60)2
[…]
*(blue!60)5*(red!60)6*(red!60)3*(red!60)1
[…]
Figure 8: The $\lambda^{1}$-slicing for $\lambda=[14,4,3,2,2,1]$. Numbers in the boxes corespond to the hook lengths after slicing.
The $ab\delta$-slicing of the diagram $\lambda=[14,4,3,2,2,1]$ can be represented as follows.
[…]
*(orange!90)2*(orange!60)1
[…]
*(orange!90)3*(orange!60)2
[…]
[δ_1]
Figure 9: The $ab\delta$-slicing for $\lambda=[14,4,3,2,2,1]$. Numbers in the boxes corespond to the hook lengths after slicing. If $E$ is the subdiagram $[3,2]$ of $\lambda$ (i.e. $E:=\{(1,1),(2,1),(3,1),(1,2),(2,2)\}$), we have $H^{*ab\delta}(\lambda,E)=1\cdot 13\cdot 12\cdot 5\cdot 1=780$.
3.2 Bounds on sliced hook products
Let us define, for any Young diagrams $\mu\subset\lambda$ and any set partition $P$ of $\lambda$, the ratio
$R_{P}(\lambda,\mu):=\frac{H(\lambda,\mu)}{H^{*P}(\lambda,\mu)}.$ (3.4)
If $P$ is the $ab\delta$-(resp. $\lambda_{1}$-, $\lambda^{1}$-)slicing, we denote the corresponding ratio by $R_{ab\delta}$ (resp. $R_{\lambda_{1}}$, $R_{\lambda^{1}}$).
We start with rewriting the ratio, in the case of a slicing with respect to the first line.
Lemma 3.4.
If $\mu,\lambda$ are two Young diagrams such that $\mu\subset\lambda$, then we have:
  • (i)
    $R_{\lambda_{1}}(\lambda,\mu)=\frac{H(\lambda,\mu_{1})}{H(\lambda_{1},\mu_{1})}$;
  • (ii)
    $R_{\lambda_{1}}(\lambda,\mu)\leq R_{\lambda_{1}}(\lambda,\lambda)$.
Proof.
Let us first prove (i). Since the first line of $\lambda$ does not appear in the hook products of boxes on the second line and above, we have $H(\lambda,\mu_{\geq 2})=H(\lambda_{\geq 2},\mu_{\geq 2})$. Therefore,
$R_{\lambda_{1}}(\lambda,\mu)=\frac{H(\lambda,\mu_{1})H(\lambda,\mu_{\geq 2})}{H(\lambda_{1},\mu_{1})H(\lambda_{\geq 2},\mu_{\geq 2})}=\frac{H(\lambda,\mu_{1})}{H(\lambda_{1},\mu_{1})}.$ (3.5)
Now we prove (ii). For each $u\in\lambda$, we have $H_{\lambda_{1}}(\lambda,u)\leq H(\lambda,u)$, i.e. $\frac{H_{\lambda_{1}}(\lambda,u)}{H(\lambda,u)}\leq 1$. Using (i), we get that
$\begin{split}\frac{R_{\lambda_{1}}(\lambda,\mu)}{R_{\lambda_{1}}(\lambda,\lambda)}&=\frac{H(\lambda,\mu_{1})/H(\lambda_{1},\mu_{1})}{H(\lambda,\lambda_{1})/H(\lambda_{1},\lambda_{1})}=\frac{H(\lambda_{1},\lambda_{1})/H(\lambda_{1},\mu_{1})}{H(\lambda,\lambda_{1})/H(\lambda,\mu_{1})}=\frac{H(\lambda_{1},\lambda_{1}\backslash\mu_{1})}{H(\lambda,\lambda_{1}\backslash\mu_{1})}\\ &=\prod_{u\in\lambda_{1}\backslash\mu_{1}}\frac{H(\lambda_{1},u)}{H(\lambda,u)}\leq 1,\end{split}$ (3.6)
where $\lambda_{1}\backslash\mu_{1}$ denotes the set of boxes that are in $\lambda_{1}$ but not in $\mu_{1}$, that is the set of boxes whose top right angle is in $\left\{(i,1){\;:\;}\mu_{1}+1\leq i\leq\lambda_{1}\right\}$. ∎
Let us now prove various bounds on $\lambda_{1}$-slicings. To ease notations, we set as before $r:=\lambda_{\geq 2}$. We start with a lemma which will be useful at several places.
Definition 3.5.
Let $p\geq 1$ and $(b_{i})_{1\leq i\leq p},(b^{\prime}_{i})_{1\leq i\leq p}$ be two (weakly) decreasing $p$-tuples of positive real numbers. We say that $(b_{1},\ldots,b_{p})\succeq(b^{\prime}_{1},\ldots,b^{\prime}_{p})$ if, for all $1\leq\ell\leq p$:
$\sum_{i=1}^{\ell}b_{i}\geq\sum_{i=1}^{\ell}b^{\prime}_{i},$ (3.7)
and
$\sum_{i=1}^{p}b_{i}=\sum_{i=1}^{p}b^{\prime}_{i}.$ (3.8)
We say that $(b_{1},\ldots,b_{p})\succ(b^{\prime}_{1},\ldots,b^{\prime}_{p})$ if, in addition, they are not equal.
Lemma 3.6.
Let $p\geq 1$ and $(a_{i})_{1\leq i\leq p},(b_{i})_{1\leq i\leq p},(b^{\prime}_{i})_{1\leq i\leq p}$ be three (weakly) decreasing $p$-tuples of positive real numbers such that $(b_{1},\ldots,b_{p})\succeq(b^{\prime}_{1},\ldots,b^{\prime}_{p})$. Then,
$\prod_{i=1}^{p}(a_{i}+b_{i})\leq\prod_{i=1}^{p}(a_{i}+b^{\prime}_{i}).$ (3.9)
Proof.
Define $f:(z_{1},\ldots,z_{p})\mapsto\prod_{i=1}^{p}(a_{i}+z_{i})$, on the simplex
$\Delta:=\left\{(x_{1},\ldots,x_{p}),x_{1}\geq\ldots\geq x_{p}>0,\sum_{i=1}^{p}x_{i}=1\right\}.$ (3.10)
We will prove that, for all $(b_{1},\ldots,b_{p})\succ(b^{\prime}_{1},\ldots,b^{\prime}_{p})$, there exists $(c_{1},\ldots,c_{p})\in\Delta$ such that $(b_{1},\ldots,b_{p})\succ(c_{1},\ldots,c_{p})\succeq(b^{\prime}_{1},\ldots,b^{\prime}_{p})$ and $f(b_{1},\ldots,b_{p})\leq f(c_{1},\ldots,c_{p})$. By continuity of $f$ (which is polynomial), this will allow us to conclude that $f(b_{1},\ldots,b_{p})\leq f(b^{\prime}_{1},\ldots,b^{\prime}_{p})$. Extending the result on simplices of the form $\Delta_{s}:=\left\{(x_{1},\ldots,x_{p}),x_{1}\geq\ldots\geq x_{p}>0,\sum_{i=1}^{p}x_{i}=s\right\}$ for $s\neq 1$ is immediate.
Fix $(b_{1},\ldots,b_{p})\succ(b^{\prime}_{1},\ldots,b^{\prime}_{p})$, and let $j:=\min\{i\in\llbracket 1,p\rrbracket,b_{i}>b^{\prime}_{i}\}$. Necessarily $j\leq p-1$. Let also $k:=\max\{i\geq j,b_{i}=b_{j}\}$ (in particular $k=j$ if $b_{j+1}<b_{j}$). By assumption, $j\leq k\leq p-1$. Let $\varepsilon\in(0,\min\{\frac{b_{k}-b^{\prime}_{k}}{2},\frac{b_{k}-b_{k+1}}{2}\})$ to be fixed later, and define $(c_{1},\ldots,c_{p})$ as: $c_{i}=b_{i}$ if $i\notin\{k,k+1\},c_{k}=b_{k}-\varepsilon,c_{k+1}=b_{k+1}+\varepsilon$. By assumption, we have $(c_{1},\ldots,c_{p})\in\Delta$ and $(b_{1},\ldots,b_{p})\succ(c_{1},\ldots,c_{p})\succeq(b^{\prime}_{1},\ldots,b^{\prime}_{p})$. In addition, we have
$\frac{f(c_{1},\ldots,c_{p})}{f(b_{1},\ldots,b_{p})}=\frac{(a_{k}+c_{k})(a_{k+1}+c_{k+1})}{(a_{k}+b_{k})(a_{k+1}+b_{k+1})}\geq 1.$ (3.11)
Indeed,
$\begin{split}&(a_{k}+c_{k})(a_{k+1}+c_{k+1})-(a_{k}+b_{k})(a_{k+1}+b_{k+1})\\ =&a_{k}(c_{k+1}-b_{k+1})+a_{k+1}(c_{k}-b_{k})+c_{k}c_{k+1}-b_{k}b_{k+1}\\ =&\varepsilon(a_{k}-a_{k+1}+b_{k}-b_{k+1})-\varepsilon^{2}>0\end{split}$ (3.12)
for $\varepsilon>0$ small enough, since $b_{k}>b_{k+1}$ by assumption. The result follows. ∎
For $k\leq n$ integers, we write
$n^{\downarrow k}:=n(n-1)...(n-k+1)=\prod_{0\leq i\leq k-1}(n-i)=\frac{n!}{(n-k)!}=k!\binom{n}{k}.$ (3.13)
Proposition 3.7.
Let $\mu\subset\lambda$ be (non-empty) Young diagrams. Then we have
  1. (a)
    We have
    $R_{\lambda_{1}}(\lambda,\mu)\leq\frac{\left(\lambda_{1}+\left\lceil\frac{u_{\leq\mu_{1}}(\lambda)}{\mu_{1}}\right\rceil\right)^{\downarrow\mu_{1}}}{\lambda_{1}^{\downarrow\mu_{1}}}.$ (3.14)
  2. (b)
    Recalling the notation $r=\lambda_{\geq 2}$, we have:
    $R_{\lambda_{1}}(\lambda,\lambda)\leq\binom{\lambda_{1}+\left\lceil\frac{r}{\lambda_{1}}\right\rceil}{\lambda_{1}}.$ (3.15)
  3. (c)
    If $r\leq\lambda_{1}$, then
    $R_{\lambda_{1}}(\lambda,\lambda)\leq 1+\frac{r}{\lambda_{1}-r+1}.$ (3.16)
  4. (d)
    Finally, we have for all $\lambda$:
    $R_{\lambda_{1}}(\lambda,\lambda)\leq e^{3\sqrt{r}}.$ (3.17)
Proof.
  1. (a)
    First observe that, as a consequence of Lemma 3.6, we have
    $\prod_{i=1}^{\mu_{1}}(\lambda_{1}-(i-1)+u_{i})\leq\prod_{i=1}^{\mu_{1}}\left(\lambda_{1}-(i-1)+\frac{u_{\leq\mu_{1}}(\lambda)}{\mu_{1}}\right).$ (3.18)
    This implies that, setting $m:=\left\lceil\frac{u_{\leq\mu_{1}}(\lambda)}{\mu_{1}}\right\rceil$, we have by Lemma 3.4 (i):
    $\begin{split}R_{\lambda_{1}}(\lambda,\mu)&=\frac{H(\lambda,\mu_{1})}{H(\lambda_{1},\mu_{1})}=\prod_{i=1}^{\mu_{1}}\frac{\lambda_{1}-(i-1)+u_{i}}{\lambda_{1}-(i-1)}\\ &\leq\prod_{i=1}^{\mu_{1}}\frac{\lambda_{1}-(i-1)+m}{\lambda_{1}-(i-1)}=\frac{(\lambda_{1}+m)^{\downarrow\mu_{1}}}{\lambda_{1}^{\downarrow\mu_{1}}}.\end{split}$ (3.19)
    This concludes the proof of (a).
  2. (b)
    In the case $\mu=\lambda$, we have $\frac{u_{\leq\mu_{1}}(\lambda)}{\mu_{1}}=\frac{r}{\lambda_{1}}$. We therefore have directly, by (a):
    $R_{\lambda_{1}}(\lambda,\lambda)\leq\frac{(\lambda_{1}+m)^{\downarrow\lambda_{1}}}{\lambda_{1}^{\downarrow\lambda_{1}}}=\binom{\lambda_{1}+m}{\lambda_{1}}.$ (3.20)
  3. (c)
    If $r=\lambda_{1}$, then (b) yields
    $R_{\lambda_{1}}(\lambda,\lambda)\leq\binom{\lambda_{1}+\left\lceil\frac{r}{\lambda_{1}}\right\rceil}{\lambda_{1}}=\lambda_{1}+1,$
    which is what we want. If $r<\lambda_{1}$, then, by Lemma 3.6, $\lambda_{1}$ and $r$ being fixed, the smallest value of $R_{\lambda_{1}}(\lambda,\lambda)$ is reached when $r$ is flat, that is, $\lambda=[\lambda_{1},n-\lambda_{1}]$. We therefore get
    $R_{\lambda_{1}}(\lambda,\lambda)\leq R_{\lambda_{1}}([\lambda_{1},n-\lambda_{1}],[\lambda_{1},n-\lambda_{1}])=\frac{\lambda_{1}+1}{\lambda_{1}-r+1}.$ (3.21)
  4. (d)
    We split the proof according to the value of $r$.
    • If $r<\lambda_{1}$, then starting from (c) we have
      $R_{\lambda_{1}}(\lambda,\lambda)\leq\frac{\lambda_{1}+1}{\lambda_{1}-r+1}\leq\frac{r+1}{r-r+1}=r+1\leq e^{\sqrt{r}},$ (3.22)
      and thus $R_{\lambda_{1}}(\lambda,\lambda)\leq e^{3\sqrt{r}}$.
    • Assume now that $r\geq\lambda_{1}$. Let $n=|\lambda|$. Note that here again $\mu=\lambda$, so that $\mu_{1}=\lambda_{1}$ and $u_{\leq\mu_{1}}=r$. Hence, $m=\left\lceil\frac{r}{\lambda_{1}}\right\rceil$, and since $r\geq\lambda_{1}$ we have $m\geq 1$. Hence, $m=\left\lceil\frac{r}{\lambda_{1}}\right\rceil\leq\frac{r}{\lambda_{1}}+1\leq 2\frac{r}{\lambda_{1}}$, so that $\lambda_{1}\leq\frac{2r}{m}$. First assume that $\lambda_{1}\leq m$. Using that $p^{\downarrow p}=p!\geq(p/e)^{p}$ for all $p\geq 0$, we therefore have, using $m\leq 2r/\lambda_{1}$ in the last inequality,
      $\binom{\lambda_{1}+m}{\lambda_{1}}=\frac{(\lambda_{1}+m)^{\downarrow\lambda_{1}}}{\lambda_{1}^{\downarrow\lambda_{1}}}\leq\frac{(2m)^{\lambda_{1}}}{(\lambda_{1}/e)^{\lambda_{1}}}=\left(\frac{2em}{\lambda_{1}}\right)^{\lambda_{1}}\leq\left(\frac{4er}{\lambda_{1}^{2}}\right)^{\lambda_{1}}.$ (3.23)
      Now observe that we can rewrite $\left(\frac{4er}{\lambda_{1}^{2}}\right)^{\lambda_{1}}=\left(\left(\frac{2\sqrt{er}}{\lambda_{1}}\right)^{\lambda_{1}}\right)^{2}$. Furthermore, for $T>0$ the function $x\in{\mathbb{R}}^{+}\mapsto(T/x)^{x}$ is maximal at $x=T/e$, we get, with $T=2\sqrt{er}$,
      $\binom{\lambda_{1}+m}{\lambda_{1}}\leq\left(e^{T/e}\right)^{2}=e^{\frac{4\sqrt{r}}{\sqrt{e}}}\leq e^{3\sqrt{r}}.$ (3.24)
      The case $m\leq\lambda_{1}$ is proved the same way by symmetry, since $\binom{\lambda_{1}+m}{\lambda_{1}}=\binom{\lambda_{1}+m}{m}$. The result follows by (b). This concludes the proof of (d).
We can now use these results to bound classical hook products and obtain a crucial inequality involving $d_{\lambda},d_{s}$ and $d_{c}$.
Proposition 3.8.
Let $\lambda$ be a (non-empty) integer partition. Then we have
  1. (a)
    We have
    $H(\lambda,\lambda)=\frac{s!}{\binom{s-1}{\lambda_{b}^{1}}}\frac{H(\lambda,\lambda_{a}^{1})}{H(\lambda_{a}^{1},\lambda_{a}^{1})}\frac{H(\lambda,\lambda_{b}^{1})}{H(\lambda_{b}^{1},\lambda_{b}^{1})}H(c,c),$ (3.25)
    and in particular
    $\frac{d_{\lambda}}{\binom{n}{c}d_{s}d_{c}}=\frac{H(\lambda_{a}^{1},\lambda_{a}^{1})}{H(\lambda,\lambda_{a}^{1})}\frac{H(\lambda_{b}^{1},\lambda_{b}^{1})}{H(\lambda,\lambda_{b}^{1})}.$ (3.26)
  2. (b)
    $\frac{d_{\lambda}}{\binom{n}{c}d_{s}d_{c}}\geq e^{-6\sqrt{c}},$ (3.27)
    where we recall that $n=|\lambda|$, $c=\lambda^{\geq 2}$, and $s=n-c=\lambda^{1}$.
Proof.
  1. (a)
    First, we have
    $\begin{split}H(\lambda,\lambda)&=sH(\lambda,\lambda_{a}^{1})H(\lambda,\lambda_{b}^{1})H(\lambda,c)\\ &=s\cdot\lambda_{a}^{1}!\cdot\lambda_{b}^{1}!\frac{H(\lambda,\lambda_{a}^{1})}{H(\lambda_{a}^{1},\lambda_{a}^{1})}\frac{H(\lambda,\lambda_{b}^{1})}{H(\lambda_{b}^{1},\lambda_{b}^{1})}H(c,c)\\ &=\frac{s!}{\binom{s-1}{\lambda_{b}^{1}}}\frac{H(\lambda,\lambda_{a}^{1})}{H(\lambda_{a}^{1},\lambda_{a}^{1})}\frac{H(\lambda,\lambda_{b}^{1})}{H(\lambda_{b}^{1},\lambda_{b}^{1})}H(c,c).\end{split}$ (3.28)
    We therefore deduce from the hook length formula and Lemma 2.3 that
    $\begin{split}\frac{d_{\lambda}}{\binom{n}{c}}=\frac{c!s!}{n!}\frac{n!}{H(\lambda,\lambda)}&=\binom{s-1}{\lambda_{b}^{1}}\frac{H(\lambda_{a}^{1},\lambda_{a}^{1})}{H(\lambda,\lambda_{a}^{1})}\frac{H(\lambda_{b}^{1},\lambda_{b}^{1})}{H(\lambda,\lambda_{b}^{1})}\frac{c!}{H(c,c)}\\ &=\binom{s-1}{\lambda_{b}^{1}}\frac{H(\lambda_{a}^{1},\lambda_{a}^{1})}{H(\lambda,\lambda_{a}^{1})}\frac{H(\lambda_{b}^{1},\lambda_{b}^{1})}{H(\lambda,\lambda_{b}^{1})}d_{c}\\ &=\frac{H(\lambda_{a}^{1},\lambda_{a}^{1})}{H(\lambda,\lambda_{a}^{1})}\frac{H(\lambda_{b}^{1},\lambda_{b}^{1})}{H(\lambda,\lambda_{b}^{1})}d_{s}d_{c},\end{split}$ (3.29)
    as desired.
  2. (b)
    By Proposition 3.7 (d) applied twice respectively to the first line and the first column of the diagrams $\lambda^{\prime}_{\geq 2}$ and $\lambda_{\geq 2}$, we get
    $\frac{H(\lambda_{a}^{1},\lambda_{a}^{1})}{H(\lambda,\lambda_{a}^{1})}\frac{H(\lambda_{b}^{1},\lambda_{b}^{1})}{H(\lambda,\lambda_{b}^{1})}\geq e^{-3\sqrt{c}}e^{-3\sqrt{c}}=e^{-6\sqrt{c}}.$ (3.30)
    Plugging this into (b) concludes the proof of (c).
3.3 Virtual degrees and augmented dimensions
If $P$ is a set partition of a diagram $\lambda\vdash n$, we can associate to it a notion of $P$-dimension via the analog of the hook length formula:
$d_{\lambda}^{*P}:=\frac{n!}{H^{*P}(\lambda,\lambda)}.$ (3.31)
The virtual degree and $ab\delta$-dimension are closely related: we have
$D(\lambda)=\frac{d_{\lambda}^{*ab\delta}}{n},$ (3.32)
where we recall that
$D(\lambda):=\frac{(n-1)!}{\prod_{i}a_{i}!b_{i}!}.$ (3.33)
Let us now define a last notion of dimension, which will prove to be very convenient in the proof of Theorem 1.5.
If $\lambda$ is a Young diagram, we define the augmented dimension $d_{\lambda}^{+}$ of $\lambda$ by
$d^{+}_{\lambda}:=\frac{n!}{H^{+}(\lambda,\lambda)},$ (3.34)
where
$H^{+}(\lambda,\lambda)=\left(\prod_{i}s_{i}\right)\left(\prod_{i}a_{i}!b_{i}!\right),$ (3.35)
as represented on the diagram below (here $s_{i}$ denotes the hook started at $\delta_{i}$ and we recall that $a_{i}=\lambda_{a}^{i}$ and $b_{i}=\lambda_{b}^{i}$).
[…]
*(orange!90)2*(orange!60)1
[…]
*(orange!90)3*(orange!60)2
[…]
[δ_1]
Note that $H^{+}(\lambda,\lambda)$ is not a sliced hook product, since we keep the full hook lengths of the boxes on the diagonal.
By definition we have
$\frac{D(\lambda)}{d_{\lambda}^{+}}=\frac{\frac{(n-1)!}{\prod_{i}a_{i}!b_{i}!}}{\frac{n!}{\left(\prod_{i}a_{i}!b_{i}!\right)\left(\prod_{i}s_{i}\right)}}=\frac{\prod_{i}s_{i}}{n}.$ (3.36)
One advantage of using $d_{\lambda}^{+}$ is the following identity (which we prove later on, in Lemma 4.7):
$d_{\lambda}^{+}=\binom{n}{c}d_{s}^{+}d_{c}^{+}.$ (3.37)
This will be convenient when bounding $d_{\lambda}^{+}$ by induction on the size of the center $c$ of $\lambda$.
4 Sharp bounds on virtual degrees
The aim of this section is to use the results of the previous sections to prove Theorem 1.5.
4.1 Strategy
In regards of (3.36), which rewrites as
$D(\lambda)=\frac{\prod_{i}s_{i}}{n}d_{\lambda}^{+},$ (4.1)
it is enough to prove the two following statements.
Proposition 4.1.
There exists a constant $C_{diag}>0$ such that, for every $n\geq 2$ and every diagram $\lambda\vdash n$, we have
$\frac{\prod_{i}s_{i}}{n}\leq d_{\lambda}^{C_{diag}/\ln n}.$ (4.2)
Proposition 4.2.
There exists a constant $C_{aug}>0$ such that, for every $n\geq 2$ and every diagram $\lambda\vdash n$, we have
$d_{\lambda}^{+}\leq d_{\lambda}^{1+C_{aug}/\ln n}.$ (4.3)
We will prove Proposition 4.1 in Section 4.2 and Proposition 4.2 in Section 4.3.
4.2 Proof of Proposition 4.1
Let us first give two general upper bounds on $\frac{\prod_{i}s_{i}}{n}$.
Lemma 4.3.
Let $n\geq 1$ and $\lambda\vdash n$ such that $c=c(\lambda)\geq 1$. Then we have
  1. (a)
    $\frac{\prod_{i}s_{i}}{n}\leq(c/\delta(c))^{\delta(c)},$ (4.4)
    where $c=\lambda^{\geq 2}$ is the center of $\lambda$ and $\delta(c):=\delta(\lambda)-1$ is the diagonal length of $c$.
  2. (b)
    Furthermore,
    $\frac{\prod_{i}s_{i}}{n}\leq e^{\sqrt{c}\ln c}.$ (4.5)
Proof.
  1. (a)
    First, since $s_{1}\leq n$, we have
    $\frac{\prod_{i}s_{i}}{n}\leq\prod_{i\geq 2}s_{i}.$ (4.6)
    Moreover, by concavity of the logarithm, we have (using that $\delta(c)=\delta(\lambda)-1$)
    $\prod_{i\geq 2}s_{i}=\exp\left(\sum_{2\leq i\leq\delta(\lambda)}\ln(s_{i})\right)\leq\exp\left(\sum_{2\leq i\leq\delta(\lambda)}\ln\left(\frac{c}{\delta(c)}\right)\right)=(c/\delta(c))^{\delta(c)},$ (4.7)
    concluding the proof of (a).
  2. (b)
    We have $\delta(c)\leq\sqrt{c}$ (since the square of side length $\delta(c)$ is included in $c$), and $s_{i}\leq c$ for each $i\geq 2$. It therefore follows that
    $\frac{\prod_{i}s_{i}}{n}\leq\prod_{i\geq 2}s_{i}\leq c^{\delta(c)}\leq c^{\sqrt{c}}=e^{\sqrt{c}\ln c},$ (4.8)
    as desired.
In Proposition 3.8 (b), we showed that $d_{\lambda}\geq\binom{s-1}{\lambda_{b}^{1}}e^{-6\sqrt{c}}\binom{n}{c}d_{c}$. Depending on the size of $c$ only some of the terms in the lower bound will be useful. We give here simple lower bounds on some of these terms.
Lemma 4.4.
Let $n\geq 1$ and $\lambda\vdash n$ such that $c\geq 1$. Then
  1. (a)
    $\binom{s-1}{\lambda_{b}^{1}}\geq s-1.$ (4.9)
  2. (b)
    $\binom{n}{c}\geq\max\left((n/c)^{c},(n/s)^{s}\right).$ (4.10)
Proof.
  1. (a)
    Since $c\geq 1$, we have $21\leq\lambda_{b}^{1}\leq s-$ 2. We therefore have
    $\binom{s-1}{\lambda_{b}^{1}}\geq\binom{s-1}{1}=s-1.$ (4.11)
  2. (b)
    This bound is a classical bound on binomial coefficients. We have
    $\binom{n}{c}=\frac{n^{\downarrow c}}{c!}\geq(n/c)^{c},$ (4.12)
    and since $\binom{n}{c}=\binom{n}{s}$, we also have $\binom{n}{c}\geq(n/s)^{s}$ by symmetry, concluding the proof.
We finally give a lower bound on $d_{\lambda}$ depending only on its diagonal length.
Lemma 4.5.
Let $n\geq 1$ and $\lambda\vdash n$. Then
$d_{\lambda}\geq(\delta/2)^{\delta(\delta-1)},$ (4.13)
where $\delta$ is the diagonal length of $\lambda$. In particular, if $n\geq(2e)^{3}$ and $\delta\geq n^{1/3}$, we have
$d_{\lambda}\geq e^{n^{2/3}/2}.$ (4.14)
Proof.
Since $d_{\lambda}$ is the number of standard tableaux of $\lambda$, and $\lambda$ contains the square $\mu:=\left[\delta^{\delta}\right]$, we have
$d_{\lambda}\geq d_{\mu}.$ (4.15)
Moreover, we claim that
$H(\mu,\mu)\leq\delta!\left((2\delta)^{\downarrow\delta}\right)^{\delta-1}.$ (4.16)
Indeed, the product of hook lengths on the top row of $\mu$ is $\delta!$. Furthermore, the hook length of a box in column $i$ is at most $12\delta-i+$ 1. Taking the product proves (4.16).
Using (4.16) and the hook length formula, we have
$\begin{split}d_{\mu}&=\frac{\delta^{2}!}{H(\mu,\mu)}\geq\frac{\delta^{2}!}{\delta!\left((2\delta)^{\downarrow\delta}\right)^{\delta-1}}=\frac{\left(\delta^{2}\right)^{\downarrow(\delta^{2}-\delta)}}{\left((2\delta)^{\downarrow\delta}\right)^{\delta-1}}=\prod_{k=2}^{\delta}\frac{(k\delta)^{\downarrow\delta}}{(2\delta)^{\downarrow\delta}}\\ &\geq\prod_{k=2}^{\delta}\frac{(k\delta)^{\delta}}{(2\delta)^{\delta}}=\frac{\left(\delta^{2}\right)^{\delta^{2}-\delta}}{\left((2\delta)^{\delta}\right)^{\delta-1}}\\ &=(\delta/2)^{\delta(\delta-1)}.\end{split}$ (4.17)
Assume now that $n\geq(2e)^{3}$ and $\delta\geq n^{1/3}$. Then we have $\delta\geq\left((2e)^{3}\right)^{1/3}=2e$ so $\delta/2\geq e$, and we also have $\delta(\delta-1)\geq\delta^{2}/2$ since $\delta\geq 2$, so $(\delta/2)^{\delta(\delta-1)}\geq e^{\delta^{2}/2}\geq e^{n^{2/3}/2}$. ∎
We now prove Proposition 4.1 for small values of $n$:
Lemma 4.6.
For any $n_{0}\geq 2$, for every $2\leq n\leq n_{0}$ and every $\lambda\vdash n$, we have
$\frac{1}{n}\prod_{i}s_{i}\leq d_{\lambda}^{C/\ln n},$ (4.18)
where $C=\frac{n_{0}(\ln n_{0})^{2}}{\ln 2}$.
Proof.
Let $n_{0}\geq 2$, $2\leq n\leq n_{0}$ and $\lambda\vdash n$. If $\lambda$ is flat (horizontal or vertical), then $d_{\lambda}=1$ and $\frac{1}{n}\prod_{i}s_{i}=1$ so the result holds. Assume now that $\lambda$ is not flat. Then $d_{\lambda}\geq 2$ so
$\frac{1}{n}\prod_{i}s_{i}\leq\prod_{1\leq i\leq\delta(\lambda)}s_{i}\leq n_{0}^{n_{0}}=2^{\frac{1}{\ln 2}n_{0}\ln n_{0}}\leq d_{\lambda}^{\frac{1}{\ln 2}n_{0}\ln n_{0}}\leq d_{\lambda}^{\frac{1}{\ln 2}(n_{0}\ln n_{0})\frac{\ln n_{0}}{\ln n}}=d_{\lambda}^{C/\ln n}.$ (4.19)
We now have all the tools to prove Proposition 4.1.
Proof of Proposition 4.1.
We prove that the result holds for $C:=C_{diag}=500$ and $n\geq e^{55}$. By Lemma 4.6, the result then extends to all $n\geq 2$, up to taking a larger constant $C_{diag}$.
We will split the proof into several cases, depending on how large $c$ is.
  • If $c=0$ then $\frac{1}{n}\prod_{i}s_{i}=\frac{n}{n}=1$ so the result holds.
  • Assume that $1\leq c\leq n^{8/9}$. First, by Proposition 3.8 (b) we have $d_{\lambda}\geq\binom{n}{c}e^{-6\sqrt{c}}$. We deduce from Lemma 4.4 (b) that
    $d_{\lambda}\geq(n/c)^{c}e^{-6\sqrt{c}}\geq n^{c/9}e^{-6\sqrt{c}}\geq n^{c/500}.$ (4.20)
    On the other hand, we have $\frac{1}{n}\prod_{i}s_{i}\leq e^{\sqrt{c}\ln c}$ by Lemma 4.3 (b). We deduce that (using that $\ln c\leq\sqrt{c}$ for all $c\geq 1$, so that $\sqrt{c}\ln c\leq c$),
    $\frac{1}{n}\prod_{i}s_{i}\leq e^{\sqrt{c}\ln c}\leq e^{c}=\left(n^{c/500}\right)^{C/\ln n}\leq d_{\lambda}^{C/\ln n}.$ (4.21)
  • Assume that $n^{8/9}\leq c\leq n-n^{2/3}$. Then
    $\frac{n}{c}\geq\frac{n}{n-n^{2/3}}=\frac{1}{1-n^{-1/3}}\geq\frac{1}{e^{-n^{-1/3}}}=e^{n^{-1/3}}$ (4.22)
    so (using again Proposition 3.8 (b) and Lemma 4.4 (b) for the first inequality)
    $d_{\lambda}\geq(n/c)^{c}e^{-6\sqrt{c}}\geq\left(e^{n^{-1/3}}\right)^{n^{8/9}}e^{-6\sqrt{c}}=e^{n^{5/9}}e^{-6\sqrt{c}}\geq e^{n^{5/9}/2}.$ (4.23)
    We conclude using Lemma 4.3 (b) (and that $c\leq n$) that
    $\frac{1}{n}\prod_{i}s_{i}\leq e^{\sqrt{c}\ln c}\leq e^{\sqrt{n}\ln n}\leq e^{500\frac{n^{5/9}}{2\ln n}}\leq d_{\lambda}^{500/\ln n}\leq d_{\lambda}^{C/\ln n}.$ (4.24)
  • Assume now that $c\geq n-n^{2/3}$, i.e. $s\leq n^{2/3}$. Then we have $\delta(\lambda)\geq n^{1/3}$ and therefore by Lemma 4.5 we have $d_{\lambda}\geq e^{n^{2/3}/2}$. We conclude, proceeding as in the previous case, that
    $\frac{1}{n}\prod_{i}s_{i}\leq e^{\sqrt{c}\ln c}\leq e^{\sqrt{n}\ln n}\leq e^{500\frac{n^{2/3}}{2\ln n}}\leq d_{\lambda}^{500/\ln n}\leq d_{\lambda}^{C/\ln n}.$ (4.25)
This concludes the proof. ∎
Remark 4.1.
In the proof of Proposition 4.1 above, (4.20) holds since $n\geq e^{5400/91}$ and (4.23) holds since $n\geq 12^{18}$. Taking $n_{0}=n^{55}$ in Lemma 4.6, we deduce that Proposition 4.1 holds (for all $n\geq 2$ if $C_{diag}\geq\max(500,\frac{n_{0}(\ln n_{0})^{2}}{\ln 2})=e^{55}(55^{2})/\ln 2$, and we can therefore take $C_{diag}=e^{64}$.
4.3 Proof of Proposition 4.2
Our aim is to show that $\frac{d_{\lambda}^{+}}{d_{\lambda}}$ is small. We will proceed by induction, slicing the external hook $s=\lambda^{1}$.
First, we recall from Proposition 3.8 (b) that the dimension of a diagram $d_{\lambda}$ and the dimension of its center $d_{c}$ are related as follows: we have $d_{\lambda}\geq e^{-6\sqrt{c}}\binom{n}{c}d_{s}d_{c}$, i.e. (recalling from Lemma 2.3 that $d_{s}=\binom{s-1}{\lambda_{a}^{1}}$):
$\frac{d_{c}}{d_{\lambda}}\leq\frac{e^{6\sqrt{c}}}{\binom{s-1}{\lambda_{a}^{1}}\binom{n}{s}}.$ (4.26)
Remark 4.2.
We also always have $d_{\lambda}\leq\binom{n}{c}d_{s}d_{c}$. Indeed, by first choosing which numbers from 1 to $n$ we place in $c$, we obtain $|\operatorname{ST}(\lambda)|\leq\binom{n}{c}|\operatorname{ST}(s)||\operatorname{ST}(c)|$. This argument was written by Diaconis and Shahshahani in their proof of cutoff for random transpositions (with $r=\lambda_{\geq 2}$ in place of $c$), see [Dia88, Fact 1 in Chapter 3D, p. 39-40] . One can therefore see Proposition 3.8 (b) as a quantified version of the intuitive approximation $d_{\lambda}\approx\binom{n}{c}d_{s}d_{c}$, for large values of $s$ and $c$.
Comparing the augmented dimensions $d_{\lambda}^{+}$ and $d_{c}^{+}$ is much simpler by design, since the hook products involved there are already sliced by definition. Also, by definition of $d_{s}^{+}$ we have $d_{s}^{+}=\frac{n!}{n\lambda_{a}^{1}!\lambda_{b}^{1}!}=\frac{s-1}{\lambda_{a}^{1}!\lambda_{b}^{1}!}=\binom{s-1}{\lambda_{a}^{1}}$, and we recall from Lemma 2.3 that $d_{s}=\binom{s-1}{\lambda_{a}^{1}}$, so that
$d_{s}^{+}=d_{s}=\binom{s-1}{\lambda_{a}^{1}}.$ (4.27)
In the next lemma we show that the aforementioned intuitive approximation for dimensions actually becomes an equality when it comes to augmented dimensions.
Lemma 4.7.
Let $n\geq 1$ and $\lambda\vdash n$. Then
$d_{\lambda}^{+}=\binom{n}{s}d_{s}^{+}d_{c}^{+}.$ (4.28)
In other words, we have
$\frac{d_{\lambda}^{+}}{d_{c}^{+}}=\binom{n}{s}{\binom{s-1}{\lambda_{a}^{1}}}.$ (4.29)
Proof.
By definition we have
$H^{+}(\lambda,\lambda)=s\lambda_{a}^{1}!\lambda_{b}^{1}!H^{+}(c,c).$ (4.30)
We therefore get
$\frac{d_{\lambda}^{+}}{d_{c}^{+}}=\frac{n!/H^{+}(\lambda,\lambda)}{c!/H^{+}(c,c)}=\frac{n^{\downarrow s}}{s\lambda_{a}^{1}!\lambda_{b}^{1}!}=\frac{n^{\downarrow s}}{s!}\frac{(s-1)!}{\lambda_{a}^{1}!\lambda_{b}^{1}!}=\binom{n}{s}{\binom{s-1}{\lambda_{a}^{1}}}.$ (4.31)
Combining (4.26) and Lemma 4.7, we obtain the following bound.
Corollary 4.8.
Let $n\geq 1$ and $\lambda\vdash n$. Then
$\frac{d_{\lambda}^{+}}{d_{\lambda}}\leq e^{6\sqrt{c}}\frac{d_{c}^{+}}{d_{c}}.$ (4.32)
We now fix the constants of Proposition 4.2. The next lemma combines the bounds which we will use just after in the proof of our result.
Lemma 4.9.
There exist constants $n_{0}\geq 3,c_{0}\geq 6$ and $C\geq 10^{3}$ such that the following hold.
  1. (i)
    For all $n\geq n_{0}$, we have
    • ($i_{1}$)
      $\sqrt{n}\leq\frac{n}{8}$,
    • ($i_{2}$)
      $n\geq 4\sqrt{n}\,\ln(n)$,
    • ($i_{3}$)
      $\frac{\ln 2}{\ln(n)}\leq\frac{1}{2}$,
    • ($i_{4}$)
      $6+\frac{6}{10^{3}}\ln(n)\leq\frac{\ln(n)}{8}$;
  2. (ii)
    for all $c\geq c_{0}$, $6+6\ln(c)\leq\frac{\sqrt{c}}{4}\ln(8/7)$;
  3. (iii)
    we have
    • ($iii_{1}$)
      $C\geq\frac{\ln(n_{0}!)\ln(n_{0})}{\ln(2)}$,
    • ($iii_{2}$)
      $C\geq 12\sqrt{c_{0}}\ln(c_{0}!)$.
Proof.
Assertions (i) and (ii) are clearly true for $c_{0},n_{0}$ large enough. Fix $c_{0},n_{0}$ so that it holds. Observe finally that, having fixed $c_{0}$ and $n_{0}$, (iii) clearly holds for $C$ large enough. ∎
For the rest of the section, we set $c_{0}$, $n_{0}$ and $C$ satisfying the conditions of Lemma 4.9.
Remark 4.3.
One can check that, for example, the values $c_{0}=10^{8},n_{0}=e^{50},C=e^{60}$ satisfy the conditions of Lemma 4.9. In other words we can take $C_{aug}=e^{60}$ in Proposition 4.2. Combining this with Remark 4.1, we deduce that in Theorem 1.5, we can take $C=e^{65}$.
We first check that our result holds for small values of $n$.
Lemma 4.10.
Let $n$ such that $2\leq n\leq n_{0}$ and let $\lambda\vdash n$. Then
$d_{\lambda}^{+}\leq d_{\lambda}^{1+C/\ln n}.$ (4.33)
Proof.
Let $2\leq n\leq n_{0}$ and $\lambda\vdash n$. If $\lambda$ is flat (horizontal or vertical), that is, $\lambda_{a}^{1}=0$ or $\lambda_{b}^{1}=0$, then $d_{\lambda}=d_{\lambda}^{+}=1$ and the result holds. Assume now that $\lambda$ is not flat. Then $d_{\lambda}\geq 2$ and
$d_{\lambda}^{+}\leq n!\leq n_{0}!=2^{\frac{\ln(n_{0}!)}{\ln 2}}\leq d_{\lambda}^{\frac{\ln(n_{0}!)}{\ln 2}}\leq d_{\lambda}^{\frac{\ln(n_{0}!)}{\ln 2}\frac{\ln n_{0}}{\ln n}}\leq d_{\lambda}^{C/\ln n}\leq d_{\lambda}^{1+C/\ln n},$ (4.34)
by Lemma 4.9 ($iii_{1}$). ∎
We now consider small values of $c$, assuming that $n$ is large enough.
Lemma 4.11.
Let $n\geq n_{0}$ and $\lambda\vdash n$ such that $c(\lambda)\leq c_{0}$. Then
$d_{\lambda}^{+}\leq d_{\lambda}^{1+C/\ln n}.$ (4.35)
Proof.
If $c=0$, we have $d_{\lambda}^{+}=d_{\lambda}$ so the result holds. Assume now that $1\leq c\leq c_{0}$. Then $\lambda_{b}^{1}\geq 1$, so that $d_{\lambda}\geq d_{s}\geq\binom{s-1}{1}=s-1\geq\sqrt{n}$ (since $n\geq n_{0}\geq 3$). Then by Corollary 4.8,
$d_{\lambda}^{+}\leq d_{\lambda}e^{6\sqrt{c}}\frac{d_{c}^{+}}{d_{c}}\leq d_{\lambda}e^{6\sqrt{c}}d_{c}^{+}\leq d_{\lambda}e^{6\sqrt{c_{0}}}c_{0}!=d_{\lambda}\sqrt{n}^{\frac{12\sqrt{c_{0}}\ln(c_{0}!)}{\ln n}}\leq d_{\lambda}^{1+C/\ln n},$ (4.36)
by Lemma 4.9 ($iii_{2}$). ∎
The next result concerns large values of $n$ and specific values of $s$.
Lemma 4.12.
Let $n\geq n_{0}$ and $s$ such that $\sqrt{n}\leq s\leq n/8$. Then
$\left(\frac{es}{n}\right)^{s/2}\leq e^{-\frac{\sqrt{n}\ln n}{8}}.$ (4.37)
Proof.
First note that since $n\geq n_{0}$, we have $\sqrt{n}\leq n/8$ by Lemma 4.9 ($i_{1}$). We define
$\begin{array}[]{lrcl}f:&\left[\sqrt{n},n/8\right]&\longrightarrow&{\mathbb{R}}\\ &x&\longmapsto&\left(\frac{ex}{n}\right)^{x/2}\end{array}$ (4.38)
and observe that $g:=\ln\circ f$ is convex (since $g^{\prime\prime}(x)=\frac{1}{2x}>0$ for $x\in\left[\sqrt{n},n/8\right]$). Hence, $g$ reaches its maximum value when $x$ is either minimal or maximal and therefore the same holds also for $f$. Moreover, $e^{2}\leq 8$ and $n\geq 4\sqrt{n}\ln n$ (by Lemma 4.9 ($i_{2}$)) so we have
$\frac{f(n/8)}{f(\sqrt{n})}=\frac{(e/8)^{n/16}}{(e/\sqrt{n})^{{\sqrt{n}/2}}}\leq\frac{\sqrt{n}^{\sqrt{n}/2}}{e^{n/16}}=\frac{n^{\sqrt{n}/4}}{e^{n/16}}\leq 1.$ (4.39)
We deduce that $f(\sqrt{n})$ is a maximum of $f$, and conclude that
$\max_{\sqrt{n}\leq s\leq n/8}f(s)\leq f(\sqrt{n})=(e/\sqrt{n})^{\sqrt{n}/2}\leq(n^{-1/4})^{\sqrt{n}/2}=e^{-\frac{\sqrt{n}\ln n}{8}},$ (4.40)
as desired. ∎
We now have all the tools to prove Proposition 4.2.
Proof of Proposition 4.2.
We proceed by (strong) induction on $n$. Our recurrence hypothesis is
$H(n):\text{\textquotedblleft}d_{\lambda}^{+}\leq d_{\lambda}^{1+C/\ln n}\text{for every }\lambda\vdash n\text{\textquotedblright}.$ ” . (4.41)
By Lemma 4.10, $H(n)$ holds for all $2\leq n\leq n_{0}$, which initializes the induction.
Let now $n\geq n_{0}$ and assume that $H(m)$ holds for every $m\leq n$. Let $\lambda\vdash n+1$.
The case $0\leq c\leq c_{0}$ follows from Lemma 4.11. Since, in addition, $s\geq\sqrt{n}$ for every diagram, we can assume from now on that $c_{0}\leq c\leq n-\sqrt{n}$.
By Corollary 4.8,
$\frac{d_{\lambda}^{+}}{d_{\lambda}^{1+\frac{C}{\ln n}}}=\frac{d_{\lambda}^{+}}{d_{\lambda}}d_{\lambda}^{-\frac{C}{\ln n}}\leq e^{6\sqrt{c}}\frac{d_{c}^{+}}{d_{c}}d_{\lambda}^{-\frac{C}{\ln n}}=e^{6\sqrt{c}}\frac{d_{c}^{+}}{d_{c}^{1+\frac{C}{\ln c}}}\left(\frac{d_{c}}{d_{\lambda}}\right)^{C/\ln c}d_{\lambda}^{\frac{C}{\ln c}-\frac{C}{\ln n}}.$ (4.42)
Applying the induction hypothesis to $c$, we deduce that
$\frac{d_{\lambda}^{+}}{d_{\lambda}^{1+\frac{C}{\ln n}}}\leq e^{6\sqrt{c}}\left(\frac{d_{c}}{d_{\lambda}}\right)^{C/\ln c}d_{\lambda}^{\frac{C}{\ln c}-\frac{C}{\ln n}}.$ (4.43)
By (4.26) and using that $d_{\lambda}\leq\binom{n}{c}d_{c}d_{s}$ (as explained in Remark 4.2), we deduce that
$\frac{d_{\lambda}^{+}}{d_{\lambda}^{1+\frac{C}{\ln n}}}$ $\leq e^{6\sqrt{c}}\left(\frac{e^{6\sqrt{c}}}{d_{s}\binom{n}{s}}\right)^{\frac{C}{\ln c}}\left(\binom{n}{c}d_{c}d_{s}\right)^{\frac{C}{\ln c}-\frac{C}{\ln n}}$
$=e^{6\sqrt{c}\left(1+\frac{C}{\ln c}\right)}\left(d_{s}\binom{n}{c}\right)^{-C/\ln n}d_{c}^{\frac{C}{\ln c}-\frac{C}{\ln n}}$
$=\left(\frac{e^{6\sqrt{c}(1+\frac{\ln c}{C})}}{\left(d_{s}\binom{n}{c}\right)^{\frac{\ln c}{\ln n}}}d_{c}^{1-\frac{\ln c}{\ln n}}\right)^{C/\ln c}$ (4.44)
$\leq\left(e^{6\sqrt{c}(1+\frac{\ln c}{C})}\binom{n}{c}^{-\frac{\ln c}{\ln n}}d_{c}^{1-\frac{\ln c}{\ln n}}\right)^{C/\ln c}$ (4.45)
To show our result, it is therefore enough to prove that
$e^{6\sqrt{c}(1+\frac{\ln c}{C})}\binom{n}{c}^{-\frac{\ln c}{\ln n}}d_{c}^{1-\frac{\ln c}{\ln n}}\leq 1.$ (4.46)
We split the proof of (4.46) depending on the value of $c$.
$n-\sqrt{n}\geq c\geq 7n/8$
In this case, we have $\sqrt{n}\leq s\leq n/8$. Then
$1-\frac{\ln c}{\ln n}=\frac{\ln n-\ln c}{\ln n}=\frac{\ln(\frac{c+s}{c})}{\ln n}=\frac{\ln(1+\frac{s}{c})}{\ln n}\leq\frac{s}{c\ln n}.$ (4.47)
In addition, by Lemma 2.3, $d_{c}\leq\sqrt{c!}\leq c^{c/2}\leq n^{c/2}$, so that
$d_{c}^{1-\frac{\ln c}{\ln n}}\leq e^{s/2}.$ (4.48)
On the other hand, since $21-s/n\geq 1/$ 2 by assumption, we have
$\frac{\ln c}{\ln n}=1+\frac{\ln(c/n)}{\ln n}=1+\frac{\ln(1-s/n)}{\ln n}\geq 1-\frac{\ln 2}{\ln n},$ (4.49)
and in particular $\frac{\ln c}{\ln n}\geq 1/2$ by Lemma 4.9 ($i_{3}$).
Recall also from Lemma 4.4 (b) that $\binom{n}{c}=\binom{n}{s}\geq(n/s)^{s}$. We therefore have, by (4.48) and Lemma 4.12,
$\left[\binom{n}{c}\right]^{-\frac{\ln c}{\ln n}}d_{c}^{1-\frac{\ln c}{\ln n}}\leq\left(\frac{s}{n}\right)^{s/2}e^{s/2}=\left(\frac{es}{n}\right)^{s/2}\leq e^{-\frac{\sqrt{n}\ln n}{8}}.$ (4.50)
We conclude, using that $\ln c\leq\ln n$ and $\sqrt{c}\leq\sqrt{n}$, that
$e^{6\sqrt{c}(1+\frac{\ln c}{C})}\binom{n}{c}^{-\frac{\ln c}{\ln n}}d_{c}^{1-\frac{\ln c}{\ln n}}\leq e^{\sqrt{n}\left(6+\frac{6\ln n}{C}-\frac{\ln n}{8}\right)}\leq 1,$ (4.51)
by Lemma 4.9 ($i_{4}$) (that we can apply since $C\geq 10^{3}$).
$c_{0}\leq c\leq 7n/8$ Since $c\geq c_{0}\geq 6$, we have by Lemma 2.3 that $d_{c}\leq\sqrt{c!}\leq(c/2)^{c/2}$. Therefore:
[…]
On the other hand we still have
$\binom{n}{c}\geq(n/c)^{c}$ (4.53)
and
$\frac{\ln c}{\ln n}=1-\frac{\ln(n/c)}{\ln n},$ (4.54)
so
$\left[\binom{n}{c}\right]^{-\frac{\ln c}{\ln n}}\leq\left(\frac{n}{c}\right)^{-c\left(1-\frac{\ln(n/c)}{\ln n}\right)},$ (4.55)
and overall we get
$\begin{split}\binom{n}{c}^{-\frac{\ln c}{\ln n}}d_{c}^{1-\frac{\ln c}{\ln n}}&\leq\left(\frac{n}{c}\right)^{-c\left(1-\frac{\ln(n/c)}{\ln n}\right)}\left(\frac{n}{c}\right)^{\frac{c}{2}\left(1-\frac{\ln(n/c)}{\ln n}\right)}2^{-\frac{c}{2}\frac{\ln(n/c)}{\ln n}}\\ &=\left(\frac{c}{n}\right)^{\frac{c}{2}\left(1-\frac{\ln(n/c)}{\ln n}\right)}2^{-\frac{c}{2}\frac{\ln(n/c)}{\ln n}}.\end{split}$ (4.56)
Observe that both factors are clearly $\leq 1$. Now, if $c\geq\sqrt{n}$ we have $21-\frac{\ln(n/c)}{\ln n}\geq 1/$ 2 so
$\left(\frac{c}{n}\right)^{\frac{c}{2}\left(1-\frac{\ln(n/c)}{\ln n}\right)}\leq\left(\frac{c}{n}\right)^{\frac{c}{4}}\leq(7/8)^{c/4}.$ (4.57)
On the other hand, if $c\leq\sqrt{n}$, we have $\frac{\ln(n/c)}{\ln n}\geq 1/2$ so
$2^{-\frac{c}{2}\frac{\ln(n/c)}{\ln n}}\leq 2^{-\frac{c}{4}}\leq(7/8)^{c/4},$ (4.58)
so we always have
$\left[\binom{n}{c}\right]^{\frac{-\ln c}{\ln n}}d_{c}^{1-\frac{\ln c}{\ln n}}\leq(7/8)^{c/4}.$ (4.59)
Hence, we get (using Lemma 4.9 (ii) and the fact that $C\geq 1$) that
$e^{6\sqrt{c}(1+\frac{\ln c}{C})}\binom{n}{c}^{-\frac{\ln c}{\ln n}}d_{c}^{1-\frac{\ln c}{\ln n}}\leq e^{6\sqrt{c}\left(1+\frac{\ln c}{C}\right)-\frac{c}{4}\ln(8/7)}\leq 1.$ (4.60)
4.4 Sharpness of the bound
We consider here two examples, namely square-shaped diagrams and almost-flat diagrams, which show sharpness of Theorem 1.5. In other words, Theorem 1.5 provides the best possible asymptotic bound that is uniform over all irreducible characters. For square-shaped diagrams $\lambda=[p^{p}]$, we show that $D(\lambda)=d_{\lambda}^{1+\frac{4\ln 2}{\ln n}(1+o(1))}$ where $n=p^{2}$, and for $\lambda=[n-2,2]$, we show that $D(\lambda)=d_{\lambda}^{1+\frac{\ln 2}{2\ln n}(1+o(1))}$. We did not optimize the value of the constant $C$ in Theorem 1.5, but we conjecture that as $n\to\infty$ we have $D(\lambda)\leq d_{\lambda}^{1+\frac{4\ln 2+o(1)}{\ln n}}$ uniformly for any $\lambda\vdash n$, and that the constant $4\ln 2$ is optimal.
4.4.1 Square-shaped diagrams
For all $n\geq 1$ such that $n$ is a perfect square (i.e. $p:=\sqrt{n}\in\mathbb{N}$), let $\lambda_{n}:=[p,\ldots,p]$ be the Young diagram of squared shape (see Fig. 10).
[…]
Figure 10: The Young diagram $\lambda_{25}$
Let us compute $d_{\lambda_{n}}$. We have
$H(\lambda_{n},\lambda_{n})=\prod_{i,j=1}^{p}(i+j-1)=\prod_{i=1}^{p}\frac{(i+p-1)!}{(i-1)!}=\frac{\prod_{j=0}^{2p-1}j!}{(\prod_{i=0}^{p-1}i!)^{2}}=\frac{G(2p+1)}{G(p+1)^{2}},$ (4.61)
where $G$ denotes Barnes’ $G$-function.
In particular, the asymptotic expansion of its logarithm is known (see e.g. [Ada14, Lemma $5.1$]): as $z\rightarrow\infty$,
$\ln G(1+z)=\frac{z^{2}}{2}\ln z-\frac{3z^{2}}{4}+O(z).$ (4.62)
This provides
$\begin{split}\ln H(\lambda_{n},\lambda_{n})&=\ln G(2p+1)-2\ln G(p+1)\\ &=\frac{(2p)^{2}}{2}\ln(2p)-\frac{3(2p)^{2}}{4}-2\frac{p^{2}}{2}\ln p+2\frac{3p^{2}}{4}+O(p)\\ &=2p^{2}\ln p+2p^{2}\ln 2-3p^{2}-p^{2}\ln p+\frac{3}{2}p^{2}+O(p)\\ &=p^{2}\ln p+(2\ln 2-\frac{3}{2})p^{2}+O(p).\end{split}$ (4.63)
Thus, we get using Stirling’s approximation:
$\begin{split}\ln d_{\lambda_{n}}&=\ln n!-\ln H(\lambda_{n},\lambda_{n})\\ &=n\ln n-n+O(\ln n)-p^{2}\ln p-(2\ln 2-\frac{3}{2})p^{2}+O(p)\\ &=2p^{2}\ln p-p^{2}-p^{2}\ln p-(2\ln 2-\frac{3}{2})p^{2}+O(p)\\ &=p^{2}\ln p+(\frac{1}{2}-2\ln 2)p^{2}+O(p).\end{split}$ (4.64)
On the other hand, we have
$\begin{split}D(\lambda_{n})=\frac{(n-1)!}{\prod a_{i}!\prod b_{i}!}=\frac{(n-1)!}{(\prod_{i=1}^{p-1}i!)^{2}}=\frac{(n-1)!}{G(1+p)^{2}}.\end{split}$ (4.65)
Using again Stirling’s formula along with (4.62), we get that
$\begin{split}\ln D(\lambda_{n})&=\ln((n-1)!)-2\ln G(p+1)\\ &=n\ln n-n+O(\ln n)-2\frac{p^{2}}{2}\ln p+2\frac{3p^{2}}{4}+O(p)\\ &=2p^{2}\ln p-p^{2}-p^{2}\ln p+\frac{3}{2}p^{2}+O(p)\\ &=p^{2}\ln p+\frac{1}{2}p^{2}+O(p).\end{split}$ (4.66)
We deduce that
$\begin{split}\ln D(\lambda_{n})-\ln d_{\lambda_{n}}=\left(2\ln 2\right)p^{2}+O(p)=\frac{4\ln 2}{\ln n}\ln d_{\lambda_{n}}(1+o(1)).\end{split}$ (4.67)
Hence, as $n\rightarrow\infty$:
$D(\lambda_{n})=d_{\lambda_{n}}^{1+\frac{4\ln 2}{\ln n}(1+o(1))}.$ (4.68)
4.4.2 Almost-flat diagrams
The second case that we consider is the case of a diagram $\mu_{n}$ of size $n$ with two lines, one of length $n-2$ and the second of length $2$ (see Fig. 11).
[…]
Figure 11: The Young diagram $\mu_{10}$
In this case, we can also compute
$H(\mu_{n},\mu_{n})=(n-4)!*(n-2)*(n-1)*2=2\frac{n!}{n(n-3)},$ (4.69)
so that
$d_{\mu_{n}}=\frac{n(n-3)}{2}.$ (4.70)
On the other hand, we have
$D(\mu_{n})=\frac{(n-1)!}{(n-3)!}=(n-1)(n-2).$ (4.71)
Hence, we have
$\begin{split}\ln\left(\frac{D(\mu_{n})}{d_{\mu_{n}}}\right)&=\ln\left(\frac{2(n-1)(n-2)}{n(n-3)}\right)\\ &=\ln 2+o(1)\\ &=\frac{\ln 2}{2\ln n}\ln d_{\mu_{n}}(1+o(1)).\end{split}$ (4.72)
Hence, we get
$D(\mu_{n})=d_{\mu_{n}}^{1+\frac{\ln 2}{2\ln n}(1+o(1))}.$ (4.73)
5 Bounds on the Witten zeta function
Recall our definition of the Witten zeta function (1.10). For $n\geq 1$, $A\subset\widehat{\mathfrak{S}_{n}}$, we denote the associated Witten zeta function by
$\zeta_{n}(A,s):=\sum_{\lambda\in A}\frac{1}{d_{\lambda}^{s}}$ (5.1)
for $s\geq 0$.
Let $\Lambda^{n}(k)=\left\{\lambda\vdash n{\;:\;}\lambda_{1}^{\prime}\leq\lambda_{1}\leq n-k\right\}$ (recall that $\lambda_{1}$ denotes the size of the first row and $\lambda_{1}^{\prime}$ the size of the first column).
Liebeck and Shalev [LS04, Proposition 2.5] proved that if $s>0$ and $k$ are fixed, then as $n\to\infty$ we have the bound
$\zeta_{n}(\Lambda^{n}(k),s)=O\left(n^{-ks}\right).$ (5.2)
This bound is very convenient. It was used for example by Larsen and Shalev [LS08] in combination with the Diaconis–Shahshahani upper bound lemma ([Dia88, Lemma 1, page 24 in Chapter 3B]), to derive mixing time results from character bounds.
However, for some applications that we will consider, we want asymptotic bounds on $\zeta_{n}(\Lambda_{k}^{n},s_{n})$ where the argument $s_{n}$ is not fixed, but tends to 00 as $n\rightarrow\infty$ instead. The goal of this section is to adapt the estimates from [LS04] to such cases.
Proposition 5.1.
Let $\alpha>0$. Let $(s_{n})_{n\geq 3}$ such that $s_{n}\geq\frac{\alpha}{\ln n}$ for all $n\geq 3$. There exist constants $k_{0}=k_{0}(\alpha)$, $n_{0}=n_{0}(\alpha)$, and $C=C(\alpha)$ such that for every $k\geq k_{0}$, for every $n\geq n_{0}$, we have
$\zeta_{n}\left(\Lambda^{n}(k),s_{n}\right)\leq Ce^{-\frac{k}{12}s_{n}\ln n}.$ (5.3)
Moreover, if $\alpha>12\ln 2$ we can take $k_{0}=1$ and $C=2$.
Proof.
We follow the proof of [LS04, Proposition $2.5$]. To ease notation we will write $\Lambda$ for $\Lambda^{(n)}(k)$. We also set $\Lambda_{1}=\Lambda_{1}^{n}(k):=\left\{\lambda\in\Lambda^{n}(k){\;:\;}\lambda_{1}\geq 2n/3\right\}$, and $\Lambda_{2}^{n}(k)=\left\{\lambda\in\Lambda^{n}(k){\;:\;}\lambda_{1}<2n/3\right\}$. we will write in the proof $\Lambda,\Lambda_{1},\Lambda_{2}$ for $\Lambda^{n}(k)$, $\Lambda_{1}^{n}(k),\Lambda_{2}^{n}(k)$ respectively. We also set $\Sigma_{i}:=\zeta_{n}(\Lambda_{i},s_{n})$ for $i\in\left\{1,2\right\}$, so that $\zeta_{n}(\Lambda,s_{n})=\Sigma_{1}+\Sigma_{2}$.
The bound obtained from [LS04] for $\Sigma_{2}:=\zeta_{n}(\Lambda_{2},s_{n})$ works ad verbum. The authors prove that there exists a constant $c>1$ such that $\Sigma_{2}\leq p(n)c^{-ns_{n}}$, where $p(n)$ is the number of partitions of the integer $n$. Since $p(n)=e^{O(\sqrt{n})}$ and $s_{n}\gtrsim 1/\ln n$, we have in particular
$\Sigma_{2}=O\left(e^{-n^{3/4}s_{n}}\right).$ (5.4)
Let us now bound $\Sigma_{1}$. We start from an intermediate bound (see [LS04, Proof of Proposition $2.5$]), namely
$\Sigma_{1}\leq\sum_{k\leq\ell\leq n/3}\frac{p(\ell)}{\binom{n-\ell}{\ell}^{s_{n}}}.$ (5.5)
From this point our proof diverges from that of [LS04]. Observe immediately that we get from (5.5) and Lemma 4.4 (b):
$\Sigma_{1}\leq\sum_{k\leq\ell\leq n/3}p(\ell)\left(\frac{\ell}{n-\ell}\right)^{\ell s_{n}}.$
We slice this sum according to whether $\ell>n^{2/3}$ or $\ell\leq n^{2/3}$. We have
$\begin{split}\Sigma_{1}\leq\Sigma_{1}^{*}+\Sigma_{1}^{**},\end{split}$ (5.6)
where $\Sigma_{1}^{*}=\sum_{k\leq\ell\leq n^{2/3}}p(\ell)\left(\frac{\ell}{n-\ell}\right)^{\ell s_{n}}$ and $\Sigma_{1}^{**}=\sum_{n^{2/3}<\ell\leq n/3}p(\ell)\left(\frac{\ell}{n-\ell}\right)^{\ell s_{n}}$.
Let us first consider $\Sigma_{1}^{**}$. Since $\ell\leq n/3$, we have $\frac{\ell}{n-\ell}\leq\frac{1}{2}$. Since $p(n)=e^{O(\sqrt{n})}$, we have $p(\ell)\leq p(n)\leq e^{n^{3/5}}$ for $n$ large enough. We therefore have for $n$ large enough, using that $s_{n}\gtrsim 1/\ln n$,
$\Sigma_{1}^{**}\leq\sum_{n^{2/3}<\ell\leq n/3}e^{n^{3/5}}\left(\frac{1}{2}\right)^{n^{2/3}s_{n}}\leq\frac{n}{3}e^{n^{3/5}}\left(\frac{1}{2}\right)^{n^{2/3}s_{n}}\leq e^{-n^{0.65}s_{n}}.$ (5.7)
We finally bound $\Sigma_{1}^{*}$. Lower bounding the denominator by $n^{5/6}$ and upper bounding $\ell$ by $n^{2/3}$, we get
$\Sigma_{1}^{*}\leq\sum_{k\leq\ell\leq n^{2/3}}p(\ell)\left(\frac{n^{2/3}}{n^{5/6}}\right)^{\ell s_{n}}=\sum_{k\leq\ell\leq n^{2/3}}p(\ell)e^{-\frac{\ell}{6}s_{n}\ln n}.$ (5.8)
Moreover since $p(\ell)=e^{O(\sqrt{\ell})}$ as $\ell\to\infty$, there exists $k_{0}=k_{0}(\alpha)$ such that, for every $\ell\geq k_{0}$, we have $p(\ell)\leq e^{\frac{\alpha}{12}\ell}$. Therefore for $k\geq k_{0}$ we have
$\Sigma_{1}^{*}$ $\leq\sum_{k\leq\ell\leq n^{2/3}}e^{\ell\left(\frac{\alpha}{12}-\frac{1}{6}s_{n}\ln n\right)}\leq\sum_{\ell=k}^{\infty}e^{\ell\left(\frac{\alpha}{12}-\frac{1}{6}s_{n}\ln n\right)}$
$=\frac{e^{k\left(\frac{\alpha}{12}-\frac{1}{6}s_{n}\ln n\right)}}{1-e^{\left(\frac{\alpha}{12}-\frac{1}{6}s_{n}\ln n\right)}}$
$\leq\frac{e^{-\frac{k}{12}s_{n}\ln n}}{1-e^{-\frac{\alpha}{12}}},$ (5.9)
using the fact that $s_{n}\ln n\geq\alpha$.
Together with (5.7) and (5.4), this ends the proof. Finally note that, if $\alpha>12\ln 2\approx 8.32$, we can simply use the bound $p(\ell)\leq 2^{\ell}\leq e^{\frac{\alpha}{12}}$ so that (5) holds with $k_{0}=1$ and $C=2$. This concludes the proof, as the bounds (5.7) and (5.4) are independent of the value of $k_{0}$. ∎
In what follows, we will write $f(n)\lll g(n)$ if $\frac{f(n)}{g(n)}\xrightarrow[n\to\infty]{}0$, that is if $f(n)=o(g(n))$.
Corollary 5.2.
Let $(s_{n})_{n\geq 3}$ be a sequence of positive real numbers. Then for every $k\geq 1$ we have
$\zeta_{n}(\Lambda^{n}(k),s_{n})\xrightarrow[n\to\infty]{}0\quad\text{ if and only if }\quad s_{n}\ggg\frac{1}{\ln n}.$ (5.10)
Proof.
First, if $s_{n}\ggg\frac{1}{\ln n}$ then $s_{n}>\frac{9}{\ln n}$ (for $n$ large enough), and, for any fixed $k\geq 1$, $\zeta_{n}(\Lambda^{n}(k),s_{n})\xrightarrow[n\to\infty]{}0$ by Proposition 5.1. Conversely, assume that $s_{n}\ggg\frac{1}{\ln n}$ does not hold. Then there exists a constant $B$ and an increasing function $\varphi{\;:\;}{\mathbb{N}}\to{\mathbb{N}}$ such that, for $n\geq 3$, $s_{\varphi(n)}\leq\frac{B}{\ln\varphi(n)}$. Fix $k\geq 1$ and, for $n$ large enough so that $n-k\geq k$, consider the diagram $[n-k,1^{k}]$. We have for all $n$, by the hook-length formula, $d_{[n-k,1^{k}]}=\binom{n-1}{k}\leq n^{k}$, so
$\zeta_{\varphi(n)}(\Lambda^{\varphi(n)}(k),s_{\varphi(n)})\geq(d_{[\varphi(n)-k,1^{k}]})^{-B/\ln\varphi(n)}\geq e^{-Bk}>0,$ (5.11)
and therefore $\zeta_{n}(\Lambda^{n}(k),s_{n})$ does not converge to 0. ∎
For $n\geq 1$ and $k\geq 0$, we denote
$\Lambda^{n}_{\mathrm{sym}}(k)=\left\{\lambda\in\widehat{\mathfrak{S}_{n}}{\;:\;}\max(\lambda_{1},\lambda^{\prime}_{1})\leq n-k\right\}.$ (5.12)
Proposition 5.3.
Let $(s_{n})_{n\geq 3}$ be a sequence of positive real numbers, and let $k\geq 1$. Then
$\zeta_{n}(\Lambda_{\mathrm{sym}}^{n}(k),s_{n})\xrightarrow[n\to\infty]{}0\quad\text{ if and only if }\quad s_{n}\ggg\frac{1}{\ln n}.$ (5.13)
In particular,
$\sum_{\lambda\in\widehat{\mathfrak{S}_{n}}^{**}}\frac{1}{d_{\lambda}^{s_{n}}}\xrightarrow[n\to\infty]{}0\quad\text{ if and only if }\quad s_{n}\ggg\frac{1}{\ln n}.$ (5.14)
Observe that the second point is exactly Proposition 1.7.
Proof.
Let $n\geq 1$. By definition we have
$\Lambda_{\mathrm{sym}}^{n}(k)=\left\{\lambda\in\widehat{\mathfrak{S}_{n}}{\;:\;}\lambda\in\Lambda^{n}(k)\text{ or }\lambda^{\prime}\in\Lambda^{n}(k)\right\}.$ (5.15)
Since the dimension of a diagram $\lambda$ is the same as the dimension of its transpose $\lambda^{\prime}$, we deduce that
$\zeta_{n}\left(\Lambda^{n}(k),s_{n}\right)\leq\zeta_{n}\left(\Lambda^{n}_{\mathrm{sym}}(k),s_{n}\right)\leq 2\zeta_{n}\left(\Lambda^{n}(k),s_{n}\right).$ (5.16)
The first statement then follows from Corollary 5.2, and the second is an application of the first one with $k=1$, since by definition $\Lambda^{n}(1)=\widehat{\mathfrak{S}_{n}}^{**}$. ∎
6 Characterization of fixed point free conjugacy classes that mix in two steps
6.1 Random walks and fixed point free conjugacy classes
Let $({\mathcal{C}}^{(n)})$ be a sequence of fixed-point free conjugacy classes (that is, such that $f_{1}({\mathcal{C}}^{(n)})=0$). We consider the sequence of random walks on ${\mathfrak{S}}_{n}$, with respective increment measures $\operatorname{\mathrm{Unif}}_{{\mathcal{C}}^{(n)}}$. Let
$\mathfrak{E}({\mathcal{C}},t):=\begin{cases}\mathfrak{S}_{n}\backslash\mathfrak{A}_{n}&\text{ if }{\mathcal{C}}\subset\mathfrak{S}_{n}\backslash\mathfrak{A}_{n}\text{ and }t\text{ is odd},\\ \mathfrak{A}_{n}&\text{ otherwise,}\end{cases}$ (6.1)
be the coset of ${\mathfrak{A}}_{n}$ on which the walk is supported after $t$ steps, and set
$\mathrm{d}^{(n)}(t):=\operatorname{d_{TV}}\left(\operatorname{\mathrm{Unif}}_{{\mathcal{C}}^{(n)}}^{*t},\operatorname{\mathrm{Unif}}_{{\mathfrak{E}}_{n}}\right)$ (6.2)
for $t\geq 0$.
Note that after 1 step the walk (started at $\operatorname{\mathrm{Id}}$) is concentrated on ${\mathcal{C}}^{(n)}$ so $\mathrm{d}^{(n)}(1)=1-o(1)$ and the walk cannot have mixed yet. Larsen and Shalev proved in [LS08] that the mixing time of such a sequence of walks is 2 or 3, i.e. (in regard of what precedes) that $\mathrm{d}^{(n)}(3)\to 0$. This settled a conjecture due to Lulov and Pak ([LP02, Conjecture 4.1]).
A natural extension is then to understand for which conjugacy classes the mixing time is 2 and for which ones it is 3 ; in other words to understand when we have $\mathrm{d}^{(n)}(2)\to 0$. The goal of this section is to prove such a characterization, Theorem 1.8.
6.2 Lower bound via a suitable observable
A way to prove a lower bound on the total variation distance between two measures $\mu$ and $\nu$ is to find a good splitting event $A$, which is an event $A$ such that $\mu(A)$ is large and $\nu(A)$ is small. For random transpositions, Diaconis and Shashahani considered the number of fixed points, which corresponds to the number of untouched cards. Here we also consider the number of fixed points but in a different way. We show that if ${\mathcal{C}}^{(n)}$ has order $n$ transpositions and $x,y\sim\operatorname{\mathrm{Unif}}_{{\mathcal{C}}^{(n)}}$ are independent uniform elements of the conjugacy class, then in the product permutation $xy$ many transpositions compensate, leading to a significantly higher probability to have many fixed points than for a uniform permutation.
Let $n\geq 1$. For $m\geq 0$ we set
$E_{m}=E_{m}^{(n)}=\left\{\sigma\in{\mathfrak{S}}_{n}{\;:\;}f_{1}(\sigma)\geq m\right\}.$ (6.3)
Let also $X=X_{n}\sim\operatorname{\mathrm{Unif}}_{{\mathfrak{A}}_{n}}$, $Y=Y_{n}\sim\operatorname{\mathrm{Unif}}_{{\mathfrak{S}}_{n}}$, and $Z=Z_{n}\sim\operatorname{Poiss}(1)$. We recall a well-known fact about the number of fixed points of random permutations.
Lemma 6.1 (Corollary 1.2 in [Nic94]).
We have
$\operatorname{d_{TV}}\left(f_{1}(Y_{n}),\operatorname{Poiss}(1)\right)\xrightarrow[n\to\infty]{}0.$ (6.4)
There are many proofs of this result and the convergence is very fast, see the discussions in [Ful24] and [DM23]. Here we will only need rough bounds on the tails for the uniform distribution on the alternating group ${\mathfrak{A}}_{n}$.
Lemma 6.2.
Let us fix $m\geq 0$. Then there exists $n_{0}=n_{0}(m)$ such that for all $n\geq n_{0}$ we have
$\operatorname{\mathrm{Unif}}_{{\mathfrak{A}}_{n}}(E_{m})\leq\frac{3}{m!}.$ (6.5)
Proof.
First, since $|{\mathfrak{S}}_{n}|=2|{\mathfrak{A}}_{n}|$, we have ${\mathbb{P}}(f_{1}(X_{n})\geq m)\leq 2{\mathbb{P}}(f_{1}(Y_{n})\geq m)$ for all $n$. Moreover, by Lemma 6.1, for $n$ large enough we have
$2{\mathbb{P}}(f_{1}(Y_{n})\geq m)\leq e{\mathbb{P}}(Z_{n}\geq m)=\sum_{j\geq m}\frac{1}{j!}\leq\frac{3}{m!}.$ (6.6)
This concludes the proof since ${\mathbb{P}}(f_{1}(X_{n})\geq m)=\operatorname{\mathrm{Unif}}_{{\mathfrak{A}}_{n}}(E_{m})$. ∎
Let us also show that the number of fixed points in subsets of $[n]$ also asymptotically behaves as a Poisson variable. We believe that this result is part of the folklore, but add a proof for completeness.
Lemma 6.3.
Let $0\leq\alpha\leq 1$. For $n\geq 1$, fix $A_{n}\subset[n]$, let $Y_{n}\sim\operatorname{\mathrm{Unif}}_{{\mathfrak{S}}_{n}}$, and define
$B_{n}(Y_{n})=\left\{i\in A_{n}{\;:\;}Y_{n}(i)=i\right\}.$ (6.7)
Assume that $|A_{n}|/n\to\alpha$ as $n\to\infty$. Then
$|B_{n}(Y_{n})|\xrightarrow[n\to\infty]{(d)}\operatorname{Poiss}(\alpha).$ (6.8)
Proof.
First, if $\alpha=1$, then a consequence of Lemma 6.1 is that, with high probability, all fixed points of $Y_{n}$ are in $A_{n}$. The result follows. The same way, if $\alpha=0$ then $\mathbb{P}(|B_{n}(Y_{n})|=0)\to 1$ as $n\rightarrow\infty$. Now suppose $10<\alpha<$ 1. Consider the set $F(Y_{n})$ of fixed points of $Y_{n}$, so that $B_{n}(Y_{n})=F(Y_{n})\cap A_{n}$. Fix two integers $j\geq k\geq 0$. We have, using the fact that $Y_{n}$ is uniform over ${\mathfrak{S}}_{n}$ and splitting over $|F(Y_{n})|$:
$\begin{split}p_{k,j}:=\mathbb{P}\left(|B_{n}(Y_{n})|=k,|F(Y_{n})|=j\right)&=\mathbb{P}(|F(Y_{n})|=j)\mathbb{P}\Big{(}|B_{n}(Y_{n})|=k\Big{|}|F(Y_{n})|=j\Big{)}\\ &=\mathbb{P}(|F(Y_{n})|=j)\frac{\binom{|A_{n}|}{k}\binom{n-|A_{n}|}{j-k}}{\binom{n}{j}}.\end{split}$ (6.9)
Moreover, as $n\to\infty$ we have $\mathbb{P}(|F(Y_{n})|=j)\to\frac{e^{-1}}{j!}$ from Lemma 6.1, and since for $r$ fixed $\binom{n}{r}\sim\frac{n^{r}}{r!}$, we also have as $n\rightarrow\infty$:
$\frac{\binom{|A_{n}|}{k}\binom{n-|A_{n}|}{j-k}}{\binom{n}{j}}\sim\frac{\frac{|A_{n}|^{k}}{k!}\frac{(n-|A_{n}|)^{j-k}}{(j-k)!}}{\frac{n^{j}}{j!}}\sim\frac{j!}{k!(j-k)!}\alpha^{k}(1-\alpha)^{j-k}.$ (6.10)
We deduce that
$p_{k,j}\xrightarrow[n\to\infty]{}\frac{e^{-1}}{k!(j-k)!}\alpha^{k}(1-\alpha)^{j-k}.$ (6.11)
Observe now that, making the change of indices $i=j-k$, we have
$\sum_{j\geq k\geq 0}\frac{e^{-1}}{k!(j-k)!}\alpha^{k}(1-\alpha)^{j-k}=e^{-1}\sum_{k\geq 0}\frac{\alpha^{k}}{k!}\sum_{i\geq 0}\frac{(1-\alpha)^{i}}{i!}=e^{-1}e^{\alpha}e^{1-\alpha}=1.$ (6.12)
Hence, there is no loss of mass and (6.11) holds uniformly for all $j\geq k\geq 0$. We deduce that for each $k\geq 0$,
$\mathbb{P}(|B_{n}(Y_{n})|=k)=\sum_{j\geq k}p_{k,j}\xrightarrow[n\to\infty]{}\sum_{j\geq k}\frac{e^{-1}}{k!(j-k)!}\alpha^{k}(1-\alpha)^{j-k}.$ (6.13)
Factoring out $\frac{e^{-1}\alpha^{k}}{k!}$ and making again the change of indices $i=j-k$, this rewrites as
$\frac{e^{-1}\alpha^{k}}{k!}\sum_{i\geq 0}\frac{(1-\alpha)^{i}}{i!}=\frac{e^{-1}\alpha^{k}}{k!}e^{1-\alpha}=\frac{e^{-\alpha}\alpha^{k}}{k!},$ (6.14)
which is the probability that a $\operatorname{Poiss}(\alpha)$ random variable is equal to $k$. This concludes the proof. ∎
Let us now show that the cancellations of transpositions after two steps leads to many fixed points. The use of unseparated pairs in the next proof is inspired from [DEG14].
Lemma 6.4.
Let $20<\alpha<1/$ 2 and $m\geq 0$. There exists $n_{1}=n_{1}(\alpha,m)$ such that for every $n\geq n_{1}$ and every conjugacy class ${\mathcal{C}}$ of $\mathfrak{S}_{n}$ such that $f_{2}({\mathcal{C}})\geq\alpha n$, we have
$\operatorname{\mathrm{Unif}}_{{\mathcal{C}}}^{*2}(E_{m})\geq\frac{e^{-\alpha}}{2}\frac{\alpha^{m/2}}{(m/2)!}$ (6.15)
Proof.
Let $n$ be a (large) integer, ${\mathcal{C}}$ be a conjugacy class of ${\mathfrak{S}}_{n}$, and $x,y\sim\operatorname{\mathrm{Unif}}_{{\mathcal{C}}}$ be independent random variables. Let $q:=2\left\lfloor\frac{\alpha n}{2}\right\rfloor$. Without loss of generality we can assume that
[…]
where $x^{\prime}$ is a permutation of the integers from $12q+$ 1 to $n$.
A way to obtain $y$ from $x$ is to consider a permutation $z\sim\operatorname{\mathrm{Unif}}_{{\mathfrak{S}}_{n}}$ and define $y=zxz^{-1}$. This corresponds to filling the cycles as follows:
$y=(z(1)\;\;z(2))(z(3)\;\;z(4))...(z(2q-1)\;\;z(2q))y^{\prime}.$ (6.17)
The transpositions of $y$ that are the same as that of $x$ then include the transpositions $(z(i)\;\;z(i+1))$ such that $z(i)$ and $z(i+1)$ are two consecutive integers that are between 1 and $2q$, a subset of which is
$S_{n,2q}(z):=\left\{i\in[2q]{\;:\;}i\text{ is odd},z(i)\in[2q],z(i)\text{ is odd, and }z(i+1)=z(i)+1\right\}.$ (6.18)
[…]
for $n$ large enough we have, letting $X\sim\operatorname{Poiss}(\alpha)$,
${\mathbb{P}}(|S_{n,2q}(z)|\geq m/2)\geq\frac{1}{2}{\mathbb{P}}(X\geq m/2)\geq\frac{1}{2}{\mathbb{P}}(X=m/2)=\frac{1}{2}e^{-\alpha}\frac{\alpha^{m/2}}{(m/2)!}.$ (6.19)
Now recall that each such fixed point of $w$ corresponds to a consecutive pair of $z$, which is a transposition of $y$ that is the same as in $x$. Hence, each element of $S_{n,2q}(z)$ corresponds to 2 fixed points in the product $xy$. We conclude that
${\mathbb{P}}(f_{1}(xy)\geq m)\geq{\mathbb{P}}(|S_{n,2q}(z)|\geq m/2)\geq\frac{e^{-\alpha}}{2}\frac{\alpha^{m/2}}{(m/2)!}.$ (6.20)
We can now use this result to prove that $E_{m}$ is a splitting event.
Proposition 6.5.
Let $20<\alpha<1/$ 2. There exists $c=c(\alpha)>0$ and $n_{2}=n_{2}(\alpha)$ such that for every $n\geq n_{2}$, and every conjugacy class ${\mathcal{C}}$ of ${\mathfrak{S}}_{n}$ such that $f_{2}({\mathcal{C}})\geq\alpha n$, we have
$\operatorname{\mathrm{Unif}}_{{\mathcal{C}}}^{*2}(E_{m})-\operatorname{\mathrm{Unif}}_{{\mathfrak{A}}_{n}}(E_{m})\geq c(\alpha),$ (6.21)
and in particular
$\operatorname{d_{TV}}\left(\operatorname{\mathrm{Unif}}_{{\mathcal{C}}^{(n)}}^{*2},\operatorname{\mathrm{Unif}}_{{\mathfrak{A}}_{n}}\right)\geq c(\alpha).$ (6.22)
Proof.
First observe that for any even integer $m\geq 2$ we have, setting $\beta=\frac{e^{-\alpha}\alpha}{2}$,
$m!\left(\frac{e^{-\alpha}}{2}\frac{\alpha^{m/2}}{(m/2)!}\right)=m^{\downarrow m/2}\frac{e^{-\alpha}}{2}\alpha^{m/2}\geq\beta^{m/2}m^{\downarrow m/2}\geq\beta^{m/2}(m/2)^{m/2}=\left(\frac{\beta m}{2}\right)^{m/2}.$ (6.23)
We now fix $m(\alpha):=\left\lceil\frac{10}{\beta}\right\rceil$. Then
$m(\alpha)!\left(\frac{e^{-\alpha}}{2}\frac{\alpha^{m(\alpha)/2}}{(m(\alpha)/2)!}\right)\geq\left(\frac{\beta m(\alpha)}{2}\right)^{m(\alpha)/2}\geq\frac{\beta m(\alpha)}{2}\geq 5.$ (6.24)
We deduce from Lemma 6.4 that for $n$ large enough
$\operatorname{\mathrm{Unif}}_{{\mathcal{C}}}^{*2}(E_{m(\alpha)})\geq\frac{5}{m(\alpha)!}.$ (6.25)
We conclude using Lemma 6.2, that for $n$ large enough
$\operatorname{\mathrm{Unif}}_{{\mathcal{C}}}^{*2}(E_{m(\alpha)})-\operatorname{\mathrm{Unif}}_{{\mathfrak{A}}_{n}}(E_{m}(\alpha))\geq\frac{5}{m(\alpha)!}-\frac{3}{m(\alpha)!}=\frac{2}{m(\alpha)!}=:c(\alpha)>0.$ (6.26)
6.3 Upper bound on characters
We now obtain an upper bound on characters applied ot fixed-point free permutations.
Lemma 6.6.
Let $n\geq 2$ and $\sigma\in{\mathfrak{S}}_{n}$ be a fixed-point free permutation. Then
$|\chi^{\lambda}(\sigma)|:=\frac{|\operatorname{\mathrm{ch}}^{\lambda}(\sigma)|}{d_{\lambda}}\leq d_{\lambda}^{-1/2}d_{\lambda}^{-\frac{1}{6}\frac{\ln g(\sigma)}{\ln n}}.$ (6.27)
where $g(\sigma)=\frac{ne^{-3C}}{\max(2f_{2}(\sigma),1)}$. In particular, we have
$d_{\lambda}|\chi^{\lambda}(\sigma)|^{2}\leq d_{\lambda}^{-\frac{1}{3}\frac{\ln g(\sigma)}{\ln n}}.$ (6.28)
Proof.
First, we have $n^{e_{2}}=\max(\Sigma_{2},1)$ i.e. $e_{2}=\frac{\ln\max(\Sigma_{2},1)}{\ln n}$, so that
$E(\sigma)\leq\frac{1}{2}e_{2}+\frac{1}{3}(1-e_{2})=\frac{1}{2}-\frac{1}{6}(1-e_{2})=\frac{1}{2}-\frac{1}{6}\frac{\ln(n/\max(\Sigma_{2},1))}{\ln n}.$ (6.29)
Therefore we have for all $C\geq 0$:
$\left(1+\frac{C}{\ln n}\right)E(\sigma)\leq\frac{1}{2}+\frac{1}{2}\frac{C}{\ln n}-\frac{1}{6}\frac{\ln(n/\max(\Sigma_{2},1))}{\ln n}=\frac{1}{2}-\frac{1}{6}\frac{\ln(ne^{-3C}/\max(\Sigma_{2},1))}{\ln n}.$ (6.30)
We deduce from Theorem 1.6 that
$|\operatorname{\mathrm{ch}}^{\lambda}(\sigma)|\leq d_{\lambda}^{\left(1+\frac{C}{\ln n}\right)E(\sigma)}\leq d_{\lambda}^{1/2}d_{\lambda}^{-\frac{1}{6}\frac{\ln\left(\frac{ne^{-3C}}{\max(\Sigma_{2},1)}\right)}{\ln n}}.$ (6.31)
This concludes the proof since $\Sigma_{2}=f_{1}+2f_{2}=2f_{2}$. ∎
Remark 6.1.
Lemma 6.6 shows in particular that if [$n\geq e^{3C}$, $f_{1}(\sigma)=0$, and $f_{2}(\sigma)\leq\frac{n}{2e^{3C}}$] then $|\operatorname{\mathrm{ch}}^{\lambda}(\sigma)|\leq\sqrt{d_{\lambda}}$.
6.4 Proof of Theorem 1.8
We finally have all the ingredients to prove Theorem 1.8.
Proof of Theorem 1.8.
First, by Proposition 6.5, if there exists some $\alpha>0$ such that $f_{2}({\mathcal{C}}^{(n)})\geq\alpha n$ (for $n$ large enough) then $\operatorname{d_{TV}}\left(\operatorname{\mathrm{Unif}}_{{\mathcal{C}}^{(n)}}^{*2},\operatorname{\mathrm{Unif}}_{{\mathfrak{A}}_{n}}\right)$ does not converge to 0. Assume now that $f_{2}({\mathcal{C}}^{(n)})=o(n)$ and let us prove that $\operatorname{d_{TV}}\left(\operatorname{\mathrm{Unif}}_{{\mathcal{C}}^{(n)}}^{*2},\operatorname{\mathrm{Unif}}_{{\mathfrak{A}}_{n}}\right)\xrightarrow[n\to\infty]{}0$.
By the Diaconis–Shahshahani upper bound lemma ([Dia88, Lemma 1 on Page 24 in Chapter 3B] but we refer here to Equation (4) in [Hou16] since we are considering the distance to the uniform measure on ${\mathfrak{A}}_{n}$), we have
$4\operatorname{d_{TV}}\left(\operatorname{\mathrm{Unif}}_{{\mathcal{C}}^{(n)}}^{*2},\operatorname{\mathrm{Unif}}_{{\mathfrak{A}}_{n}}\right)^{2}\leq\sum_{\lambda\in\widehat{\mathfrak{S}_{n}}^{**}}\left(d_{\lambda}|\chi^{\lambda}({\mathcal{C}}^{(n)})|^{2}\right)^{2}.$ (6.32)
Set $g({\mathcal{C}}^{(n)})=\frac{ne^{-3C}}{\max(2f_{2}({\mathcal{C}}^{(n)}),1)}$ as in Lemma 6.6 and set also $s_{n}({\mathcal{C}}^{(n)}):=\frac{2}{3}\frac{\ln g({\mathcal{C}}^{(n)})}{\ln n}$.
Then applying Lemma 6.6 to the right hand side of (6.32) we obtain
$4\operatorname{d_{TV}}\left(\operatorname{\mathrm{Unif}}_{{\mathcal{C}}^{(n)}}^{*2},\operatorname{\mathrm{Unif}}_{{\mathfrak{A}}_{n}}\right)^{2}\leq\sum_{\lambda\in\widehat{\mathfrak{S}_{n}}^{**}}\left(d_{\lambda}^{-\frac{1}{3}\frac{\ln g(\sigma)}{\ln n}}\right)^{2}=\sum_{\lambda\in\widehat{\mathfrak{S}_{n}}^{**}}\frac{1}{(d_{\lambda})^{s_{n}({\mathcal{C}}^{(n)})}}.$ (6.33)
Finally, we have $f_{2}({\mathcal{C}}^{(n)})=o(n)$ by assumption, which is equivalent to $s_{n}({\mathcal{C}}^{(n)})\ggg\frac{1}{\ln n}$. It then follows from Proposition 5.3 that
$\operatorname{d_{TV}}\left(\operatorname{\mathrm{Unif}}_{{\mathcal{C}}^{(n)}}^{*2},\operatorname{\mathrm{Unif}}_{{\mathfrak{A}}_{n}}\right)\xrightarrow[n\to\infty]{}0,$ (6.34)
which concludes the proof of Theorem 1.8. ∎
Acknowledgements
We thank Nathanaël Berestycki, Persi Diaconis, Valentin Féray, and Sam Olesker-Taylor for helpful conversations and comments.
L.T. was supported by the Pacific Institute for the Mathematical Sciences and the Simons fundation. P.T. was supported by the Austrian Science Fund (FWF) under grant 10.55776/P33083.